### v2

# https ://ydeng.perso.math.cnrs.fr

[…]
https ://ydeng.perso.math.cnrs.fr
[…]
https ://sites.google.com/site/yamanoimath/
Linear Shafarevich Conjecture in positive characteristic, Hyperbolicity and Applications
Ya Deng ya.deng@math.cnrs.fr, deng@imj-prg.fr CNRS, Institut Élie Cartan de Lorraine, Université de Lorraine, 54506, Nancy & Institut de Mathématiques de Jussieu-Paris Rive Gauche, Sorbonne Université, Campus Pierre et Marie Curie, 4 place Jussieu, 75252 Paris Cedex 05, France    Katsutoshi Yamanoi yamanoi@math.sci.osaka-u.ac.jp Department of Mathematics, Graduate School of Science,Osaka University, Toyonaka, Osaka 560-0043, Japan
Résumé
Given a complex quasi-projective normal variety $X$ and a linear representation $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ with $K$ any field of positive characteristic, we mainly establish the following results :
  1. (a)
    the construction of the Shafarevich morphism ${\rm sh}_{\varrho}:X\to{\rm Sh}_{\varrho}(X)$ associated with $\varrho$.
  2. (b)
    In cases where $X$ is projective, $\varrho$ is faithful and the $\Gamma$-dimension of $X$ is at most two (e.g. $\dim X=2$), we prove that the Shafarevich conjecture holds for $X$ : the universal covering of $X$ is holomorphically convex.
  3. (c)
    In cases where $\varrho$ is big, we prove that the Green-Griffiths-Lang conjecture holds for $X$ : $X$ is of log general type if and only it is pseudo Picard or Brody hyperbolic.
  4. (d)
    When $\varrho$ is big and the Zariski closure of $\varrho(\pi_{1}(X))$ is a semisimple algebraic group, we prove that $X$ is pseudo Picard hyperbolic, and strongly of log general type.
  5. (e)
    If $X$ is special or $h$-special, then $\varrho(\pi_{1}(X))$ is virtually abelian.
We also prove Claudon-Höring-Kollár’s conjecture for complex projective manifolds with linear fundamental groups of any characteristic.
keywords:
Shafarevich conjecture, holomorphic convexity, Shafarevich morphism, Green-Griffiths-Lang conjecture, special loci, pseudo Picard hyperbolicity, compatifiable universal covering, Campana’s special varieties, Campana’s abelianity conjecture
0 Main results
0.1 Existence of the Shafarevich morphism
The Shafarevich conjecture stipulates that the universal covering of a complex projective variety is holomorphically convex. If this conjecture holds true, it implies the existence of the Shafarevich morphism. Over the past three decades, this conjecture has been extensively studied when considering cases where fundamental groups are subgroups of complex general linear groups, referred to as the linear Shafarevich conjecture. Drawing upon the robust techniques of non-abelian Hodge theories established by Simpson [Sim88, Sim92] and Gromov-Schoen [GS92], linear Shafarevich conjecture has been studied in [Kat97, KR98, Eys04, EKPR12, CCE15, DYK23], to quote only a few. It is natural to ask whether the conjecture holds when the fundamental groups of algebraic varieties are subgroups of general linear groups in positive characteristic. In this paper we address this question along with the exploration of hyperbolicity and algebro-geometric properties of these algebraic variety. The first result of this paper is the construction of the Shafarevich morphism.
Theorem A (=Theorem˜2.9).
Let $X$ be a quasi-projective normal variety and
$\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$
be a linear representation, where $K$ is a field of positive characteristic. Then there exists a dominant (algebraic) morphism ${\rm sh}_{\varrho}:X\to{\rm Sh}_{\varrho}(X)$ over a quasi-projective normal variety ${\rm Sh}_{\varrho}(X)$ with connected general fibers such that for any connected Zariski closed subset $Z\subset X$, the following properties are equivalent :
  1. (a)
    ${\rm sh}_{\varrho}(Z)$ is a point ;
  2. (b)
    $\varrho({\rm Im}[\pi_{1}(Z)\to\pi_{1}(X)])$ is finite ;
  3. (c)
    for each irreducible component $Z_{o}$ of $Z$, $\varrho^{ss}({\rm Im}[\pi_{1}(Z_{o}^{\rm norm})\to\pi_{1}(X)])$ is finite, where $\varrho^{ss}:\pi_{1}(X)\to{\rm GL}_{N}(\bar{K})$ is the semisimplification of $\varrho$ and $Z_{o}^{\rm norm}$ denotes the normalization of $Z_{o}$.
The above morphism ${\rm sh}_{\varrho}:X\to{\rm Sh}_{\varrho}(X)$ will be called the Shafarevich morphism associated with $\varrho$. We remark that in our previous work [DYK23], Theorem˜A was proved when ${\rm char}\,K=0$ and $\varrho$ is semisimple, with a weaker statement that ${\rm sh}_{\varrho}:X\to{\rm Sh}_{\varrho}(X)$ is algebraic in the function field level. See also the paper by Brunebarbe [Bru23], which independently obtained a similar result.
We also prove the following theorem on the Shafarevich conjecture.
Theorem B (=Theorem˜6.23).
Let $X$ be a projective normal variety and let
$\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$
be a faithful representation where $K$ is a field of positive characteristic. If the $\Gamma$-dimension (see Definition˜6.22) of $X$ is at most two (e.g. when $\dim X\leq 2$), then the universal covering $\widetilde{X}$ of $X$ is holomorphically convex.
0.2 On the Green-Griffiths-Lang conjecture
Building on the methods utilized in establishing Theorem˜A, together with the techniques developed in [CDY22], we prove the following theorem on the generalized Green-Griffiths-Lang conjecture. A stronger and more refined result will be stated in Theorem˜D.
[…]
Let $X$ be a complex quasi-projective normal variety. Let $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be a big representation where $K$ is a field of positive characteristic. Then the following properties are equivalent :
  1. (a)
    $X$ is of log general type ;
  2. (b)
    $X$ is strongly of log general type ;
  3. (c)
    $X$ is pseudo Picard hyperbolic, that is, there exists a proper Zariski closed subset $\Xi\subsetneq X$ such that any holomorphic map $f:\mathbb{D}^{*}\to X$ from the punctured disk with essential singularity at the origin has image $f(\mathbb{D}^{*})\subset\Xi$.
  4. (d)
    $X$ is pseudo Brody hyperbolic, that is, there exists a proper Zariski closed subset $\Xi\subsetneq X$ such that any non-constant holomorphic map $f:\mathbb{C}\to X$ has image $f(\mathbb{C})\subset\Xi$.
We say a quasi-projective variety $X$ is strongly of log general type if there exists a proper Zariski closed subset $\Xi\subsetneq X$ such that any positive dimensional closed subvariety $V\subset X$ is of log general type provided that $V\not\subset\Xi$. Recall that a representation $\varrho:\pi_{1}(X)\to G(K)$ is said to be big, (or generically large in [Kol95]), if for any closed irreducible subvariety $Z\subset X$ containing a very general point of $X$, $\varrho({\rm Im}[\pi_{1}(Z^{\rm norm})\to\pi_{1}(X)])$ is infinite. It is worth noticing that a stronger notion exists : a representation $\varrho$ is called large if $\varrho({\rm Im}[\pi_{1}(Z^{\rm norm})\to\pi_{1}(X)])$ is infinite for any closed subvariety $Z$ of $X$. We remark that in [CDY22] we prove Theorem˜C when ${\rm char}\,K=0$ and $\varrho$ is semisimple. It is worthwhile to mention that in the case ${\rm char}\,K>0$, $\varrho$ is not required to be semisimple.
We would like to refine Theorem˜C to compare the non-hyperbolicity locus of the hyperbolicity notions in Theorem˜C. We first introduce a notion of special loci ${\rm Sp}(\varrho)$ for any big representation $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ which measures the "non-large locus" of $\varrho$.
Definition 0.1.
Let $X$ be a smooth quasi-projective variety. Let $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be a representation where $K$ is a field. We define
${\rm Sp}(\varrho):=\overline{\bigcup_{\iota:Z\hookrightarrow X}Z}^{\rm Zar},$
where $\iota:Z\hookrightarrow X$ ranges over all positive dimensional closed subvarieties of $X$ such that $\iota^{*}\varrho(\pi_{1}(Z))$ is finite.
Additionally, as in [CDY22, Definition 0.1], we can introduce four special subsets $\mathrm{Sp}_{\mathrm{sab}}$, $\mathrm{Sp}_{\mathrm{alg}}(X)$, $\mathrm{Sp}_{\mathrm{h}}(X)$ and $\mathrm{Sp}_{\mathrm{p}}(X)$ of $X$ that measure the non-hyperbolicity locus of the hyperbolicity notions in Theorem˜C from different perspectives.
Definition 0.2.
Let $X$ be a quasi-projective normal variety. We define
  1. (i)
    $\mathrm{Sp}_{\mathrm{sab}}(X):=\overline{\bigcup_{f}f(A_{0})}^{\mathrm{Zar}}$, where $f$ ranges over all non-constant rational maps
    $f:A\dashrightarrow X$
    from all semi-abelian varieties $A$ to $X$ such that $f$ is regular on a Zariski open subset $A_{0}\subset A$ whose complement $A\backslash A_{0}$ has codimension at least two ;
  2. (ii)
    $\mathrm{Sp}_{\mathrm{h}}(X):=\overline{\bigcup_{f}f(\mathbb{C})}^{\mathrm{Zar}}$, where $f$ ranges over all non-constant holomorphic maps from $\mathbb{C}$ to $X$ ;
  3. (iii)
    $\mathrm{Sp}_{\mathrm{alg}}(X):=\overline{\bigcup_{V}V}^{\mathrm{Zar}}$, where $V$ ranges over all positive-dimensional closed subvarieties of $X$ which are not of log general type ;
  4. (iv)
    $\mathrm{Sp}_{\mathrm{p}}(X):=\overline{\bigcup_{f}f(\mathbb{D}^{*})}^{\mathrm{Zar}}$, where $f$ ranges over all holomorphic maps from the punctured disk $\mathbb{D}^{*}$ to $X$ with essential singularity at the origin, i.e., $f$ has no holomorphic extension $\bar{f}:\mathbb{D}\to\overline{X}$ to a projective compactification $\overline{X}$.
Subsequently, we establish a theorem concerning these special subsets, thereby refining Theorem˜C.
Theorem D (=Lemmas˜3.4 and 3.6).
Let $X$ be a quasi-projective normal variety. Let $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be a big representation where $K$ is a field of positive characteristic. Then ${\rm Sp}(\varrho)$ is a proper Zariski closed subset of $X$, and we have
$\mathrm{Sp}_{\mathrm{sab}}(X)\setminus{\rm Sp}(\varrho)=\mathrm{Sp}_{\mathrm{alg}}(X)\setminus{\rm Sp}(\varrho)=\mathrm{Sp}_{\mathrm{p}}(X)\setminus{\rm Sp}(\varrho)=\mathrm{Sp}_{\mathrm{h}}(X)\setminus{\rm Sp}(\varrho).$
We have ${\rm Sp}_{\bullet}(X)\subsetneq X$ if and only if $X$ is of log general type, where $\rm Sp$$_{\bullet}$ denotes any of $\mathrm{Sp}_{\mathrm{sab}}$, $\mathrm{Sp}_{\mathrm{alg}}$, $\mathrm{Sp}_{\mathrm{h}}$ or $\mathrm{Sp}_{\mathrm{p}}$.
0.3 How fundamental groups determine hyperbolicity
It is natural to explore how the fundamental groups of algebraic varieties determine their hyperbolicity properties. In our previous work [CDY22], we provided a characterization based on representations of fundamental groups into complex general linear groups. In this paper, we establish analogous results concerning representations in positive characteristic fields.
Theorem E (=Theorem˜3.7).
Let $X$ be a quasi-projective normal variety. Let
$\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$
be a big representation where $K$ is an algebraically closed field of positive characteristic. If the Zariski closure $\varrho(\pi_{1}(X))$ is a semisimple algebraic group over $K$, then ${\rm Sp}_{\bullet}(X)\subsetneq X$, where $\rm Sp$$_{\bullet}$ denotes any of $\mathrm{Sp}_{\mathrm{sab}}$, $\mathrm{Sp}_{\mathrm{alg}}$, $\mathrm{Sp}_{\mathrm{h}}$ or $\mathrm{Sp}_{\mathrm{p}}$, in particular, $X$ is of log general type.
It is worthwhile mentioning that when ${\rm char}\,K=0$, Theorem˜E was proved in [CDY22, Theorem A]. We also remark that the condition in Theorem˜3.7 is sharp (cf. Remark˜3.9).
0.4 Some applications
Theorem˜E has various applications. Campana’s abelianity conjecture [Cam04] predicts that a smooth projective variety $X$ that is special has a virtually abelian fundamental group. Our first application of Theorem˜E is the proof of this conjecture in the context of representations in positive characteristic.
Theorem F (=Theorem˜4.12).
Let $X$ be a smooth quasi-projective variety, and let $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be any representation where $K$ is a field of positive characteristic. If $X$ is special or $h$-special (cf. Definitions˜4.1 and 4.2), then $\varrho(\pi_{1}(X))$ is virtually abelian.
Note that in cases when $X$ is projective and ${\rm char}\,K=0$, Theorem˜F was proved by Campana [Cam04] (for $X$ special) and the second author [Yam10] (for $X$ Brody special). It is important to mention that Theorem˜F does not hold when ${\rm char}\,K=0$ as in [CDY22, Example 11.26] we constructed a special and Brody special smooth quasi-projective variety with nilpotent fundamental group but not virtually abelian. In cases where ${\rm char}\,K=0$, in [CDY22] we prove that $\varrho(\pi_{1}(X))$ is virtually nilpotent (cf. Theorem˜4.7).
As a first consequence of Theorem˜F, we can address a conjecture proposed by Claudon, Höring, and Kollár concerning algebraic varieties with compactifiable universal coverings (cf. Conjecture˜6.1).
Corollary G (=Corollary˜6.2).
Let $X$ be a smooth projective variety with an infinite fundamental group $\pi_{1}(X)$, such that its universal covering $\widetilde{X}$ is a Zariski open subset of some compact Kähler manifold. If there exists a faithful representation $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K),$ where $K$ is any field of any characteristic, then the Albanese map of $X$ is (up to finite étale cover) locally isotrivial with simply connected fiber $F$. In particular we have $\widetilde{X}\simeq F\times\mathbb{C}^{q(X)}$ with $q(X)$ the irregularity of $X$.
As another application of Theorem˜F, we provide a characterization of semiabelian variety.
Corollary H (=Proposition˜4.13).
Let $X$ be a smooth quasi-projective variety, and let $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be a big representation where $K$ is a field of positive characteristic.
  1. (a)
    If $X$ is special or $h$-special, then after replacing $X$ by some finite étale cover, its quasi-Albanese map $\alpha:X\to A$ is birational and $\alpha_{*}:\pi_{1}(X)\to\pi_{1}(A)$ is an isomorphism.
  2. (b)
[…]
It is worth mentioning that in [CDY22] we proved Corollary˜H in cases where ${\rm char}\,K=0$ and $\varrho$ is big and reductive.
Lastly, we apply Theorem˜E to obtain a structure theorem for quasi-projective varieties $X$ for which there exists a big representation $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ where $K$ is a field of positive characteristic. See Theorem˜5.2.
0.5 Structure of the paper
The paper presents several results from different perspectives. For the readers’ convenience, we list in Figure 1 the relationships between main theorems.
Theorem˜ATheorem˜BTheorem˜CTheorem˜DTheorem˜ETheorem˜5.2Theorem˜FCorollary˜HCorollary˜G
Figure 1: Relationships between Main Theorems
We remark that Theorems˜A, B, C, D and E are entirely novel results, even in cases where $X$ is a projective variety. Their proofs differ from those used in studying complex reductive representations of fundamental groups. In a forthcoming work, we will extend Theorem˜B to arbitrary projective normal varieties.
Convention and notation
In this paper, we use the following conventions and notations :
  • Quasi-projective varieties and their closed subvarieties are assumed to be positive-dimensional and irreducible unless specifically mentioned otherwise. Zariski closed subsets, however, may be reducible.
  • Fundamental groups are always referred to as topological fundamental groups.
  • If $X$ is a complex space, its normalization is denoted by $X^{\mathrm{norm}}$.
  • $\mathbb{D}$ denotes the unit disk in $\mathbb{C}$, and $\mathbb{D}^{*}$ denotes the punctured unit disk.
  • For an algebraic group $G$, we denote by $\mathcal{D}G$ its derived group.
  • For any prime number $p$, we denote by ${\rm GL}(N,\mathbb{F}_{p})$ the general linear group over $\mathbb{F}_{p}$. If $K$ is a field with ${\rm char}\,K=p$, we denote by ${\rm GL}_{N}(K)$ its $K$-points.
  • For a finitely generated group $\Gamma$, any field $K$ and any representation $\varrho:\Gamma\to{\rm GL}_{N}(K)$, we denote by $\varrho^{ss}:\Gamma\to{\rm GL}_{N}(\bar{K})$ the semisimplification of $\varrho$, where $\bar{K}$ denotes the algebraic closure of $K$. The representation $\varrho$ is reductive if the Zariski closure of $\varrho(\Gamma)$ is a reductive group. We note that $\varrho^{ss}$ is a semisimple representation, thus reductive (cf. [Mil17, Corollary 19.18]).
1 Technical preliminary
1.1 Katzarkov-Eyssidieux reduction
Definition 1.1 (Katzarkov-Eyssidieux reduction).
Let $X$ be a complex smooth quasi-projective variety, and let $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be a linear representation where $K$ is a non-archimedean local field. A morphism $s_{\varrho}:X\to S_{\varrho}$ to a complex normal quasi-projective variety $S_{\varrho}$ is called Katzarkov-Eyssidieux reduction map if
  • $s_{\varrho}$ is dominant and has connected general fibers, and
  • for every connected Zariski closed subset $T$ of $X$, the image $s_{\varrho}(T)$ is a point if and only if the image $\varrho({\rm Im}[\pi_{1}(T)\to\pi_{1}(X)])$ is a bounded subgroup of $\mathrm{GL}_{N}(K)$.
Recall that a subgroup of $\mathrm{GL}_{N}(K)$ is said to be bounded if it is contained in a maximal compact subgroup of $\mathrm{GL}_{N}(K)$. Note that a maximal compact subgroup of $\mathrm{GL}_{N}(K)$ is, up to conjugation, equal to $\mathrm{GL}_{N}(\mathcal{O}_{K})$.
When $X$ is projective, we may easily see that $s_{\varrho}:X\to S_{\varrho}$ is unique up to isomorphism, if it exists. In our previous work [CDY22] jointly with Cadorel, we establish the existence of Katzarkov-Eyssidieux reduction map for reductive representations. This generalized previous work by Katzarkov [Kat97] and Eyssidieux [Eys04] from projective varieties to the quasi-projective cases. Here we state a stronger result, which is implicitly contained in our paper [DYK23].
Theorem 1.2.
Let $X$ be a complex smooth quasi-projective variety, and let
$\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$
be a linear representation where $K$ is a non-archimedean local field. Then there exists a quasi-projective normal variety $S_{\varrho}$ and a dominant morphism $s_{\varrho}:X\to S_{\varrho}$ with connected general fibers, such that for any connected Zariski closed subset $T$ of $X$, the following properties are equivalent :
  1. (a)
    the image $\varrho({\rm Im}[\pi_{1}(T)\to\pi_{1}(X)])$ is a bounded subgroup of ${\rm GL}_{N}(K)$.
  2. (b)
    For every irreducible component $T_{o}$ of $T$, the image $\varrho({\rm Im}[\pi_{1}(T_{o}^{\rm norm})\to\pi_{1}(X)])$ is a bounded subgroup of ${\rm GL}_{N}(K)$.
  3. (c)
    The image $s_{\varrho}(T)$ is a point.∎
Démonstration.
Let $\varrho^{ss}:\pi_{1}(X)\to{\rm GL}_{N}(\bar{K})$ be the semisimplification of $\varrho$. We assume that $L/K$ is a finite extension such that $\varrho^{ss}:\pi_{1}(X)\to{\rm GL}_{N}(L)$. It is proven in [CDY22, Theorem H] that the Katzarkov-Eyssidieux reduction map $s_{\varrho^{ss}}:X\to S_{\varrho^{ss}}$ for $\varrho^{ss}$ exists and satisfies the properties in the theorem.
On the other hand, we have the following result.
Claim 1.3 ([DYK23, Lemma 3.7]).
Let $K$ be a non-archimedean local field and $\Gamma$ be a finitely generated group. If $\{\varrho_{i}:\Gamma\to{\rm GL}_{N}(\bar{K})\}_{i=1,2}$ are two linear representations such that there semisimplifications are conjugate, then $\varrho_{1}$ is bounded if and only if $\varrho_{2}$ is bounded. ∎
Therefore, if we define $s_{\varrho}$ to be $s_{\varrho^{ss}}:X\to S_{\varrho^{ss}}$, it satisfies the properties required in the theorem. Indeed, let $T\subset X$ be a connected Zariski closed subset. Set $\Gamma=\pi_{1}(T)$. Then $\Gamma$ is finitely generated. Let $\iota:\Gamma\to\pi_{1}(X)$ be a natural morphism. Note that the semisimplifications of two composite representations $\varrho\circ\iota:\Gamma\to{\rm GL}_{N}(\bar{K})$ and $\varrho^{ss}\circ\iota:\Gamma\to{\rm GL}_{N}(\bar{K})$ are conjugate. Hence by Claim˜1.3, $\varrho\circ\iota$ is bounded iff $\varrho^{ss}\circ\iota$ is bounded. Hence the image $s_{\varrho^{ss}}(T)$ is a point iff $\varrho({\rm Im}[\pi_{1}(T)\to\pi_{1}(X)])$ is a bounded subgroup of ${\rm GL}_{N}(K)$. Similarly $s_{\varrho^{ss}}(T)$ is a point iff $\varrho({\rm Im}[\pi_{1}(T_{o}^{\rm norm})\to\pi_{1}(X)])$ is a bounded subgroup of ${\rm GL}_{N}(K)$ for every irreducible component $T_{o}$ of $T$. Thus $s_{\varrho^{ss}}:X\to S_{\varrho^{ss}}$ satisfies the properties required in the theorem. ∎
According to this theorem, the two properties (a) and (b) are equivalent for every connected Zariski closed subset $T$ of $X$. Hence for every Katzarkov-Eyssidieux reduction map $s_{\varrho}:X\to S_{\varrho}$, the three statements (a), (b) and (c) are equivalent.
Remark 1.4.
The proof of Theorem˜1.2 shows that for every linear representation
$\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K),$
where $K$ is a non-archimedean local field, the Katzarkov-Eyssidieux reduction map
$s_{\varrho^{ss}}:X\to S_{\varrho^{ss}}$
for the semisimplification $\varrho^{ss}$ is the Katzarkov-Eyssidieux reduction map for $\varrho$.
The following lemma proved in [DYK23] will be used throughout this paper.
Lemma 1.5 ([DYK23, Lemma 1.28]).
Let $V$ be a quasi-projective normal variety and let $(f_{\lambda}:V\to S_{\lambda})_{\lambda\in\Lambda}$ be a family of morphisms into quasi-projective varieties $S_{\lambda}$. Then there exist a quasi-projective normal variety $S_{\infty}$ and a morphism $f_{\infty}:V\to S_{\infty}$ such that
  • $f_{\infty}$ is dominant and has connected general fibers,
  • for every subvariety $Z\subset V$, $f_{\infty}(Z)$ is a point if and only if $f_{\lambda}(Z)$ is a point for every $\lambda\in\Lambda$, and
  • there exist $\lambda_{1},\ldots,\lambda_{n}\in\Lambda$ such that $f_{\infty}:V\to S_{\infty}$ is the quasi-Stein factorization of $(f_{1},\ldots,f_{n}):V\to S_{\lambda_{1}}\times\cdots\times S_{\lambda_{n}}$. ∎
Such $f_{\infty}:V\to S_{\infty}$ is called the simultaneous Stein factorization of $(f_{\lambda}:V\to S_{\lambda})_{\lambda\in\Lambda}$.
Lemma 1.6.
Let $p$ be a prime number. Let $\Lambda$ be a non-empty set of reductive representations $\tau:\pi_{1}(X)\to\mathrm{GL}_{N_{\tau}}(K_{\tau})$, where $K_{\tau}$ are local fields of $\mathrm{char}\ K_{\tau}=p$. Then there exist
  • a local field $K$ of $\mathrm{char}\ K=p$,
  • a positive integer $N>0$,
  • a reductive representation $\varrho:\pi_{1}(X)\to\mathrm{GL}_{N}(K)$
such that the simultaneous Stein factorization of Katzarkov-Eyssidieux reduction maps
$(s_{\tau}:X\to S_{\tau})_{\tau\in\Lambda}$
coincides with Katzarkov-Eyssidieux reduction map of $\varrho$. Moreover we have
$\cap_{\tau\in\Lambda}\mathrm{ker}(\tau^{ss})\subset\mathrm{ker}(\varrho).$
Démonstration.
Let $\sigma:X\to\Sigma$ be the simultaneous Stein factorization of $(s_{\tau}:X\to S_{\tau})_{\tau\in\Lambda}$. Then $\Sigma$ is normal, and $\sigma:X\to\Sigma$ is dominant and has connected general fibers. By Lemma˜1.5, there exist $\tau_{1},\ldots,\tau_{n}\in\Lambda$ such that $\sigma:X\to\Sigma$ is the quasi-Stein factorization of $(s_{\tau_{1}},\ldots,s_{\tau_{n}}):X\to S_{\tau_{1}}\times\cdots\times S_{\tau_{n}}$. Note that each $K_{\tau_{i}}$ is a finite extension of the Laurent series $\mathbb{F}_{p}((t))$ over the finite field $\mathbb{F}_{p}$. We may take a local field $K$ of $\mathrm{char}\ K=p$ such that $K_{\tau_{i}}\subset K$ and $\tau_{i}^{ss}:\pi_{1}(X)\to\mathrm{GL}_{N}(K)$ for all $i=1,\ldots,n$. Set $N=N_{\tau_{1}}+\cdots+N_{\tau_{n}}$. We define $\varrho:\pi_{1}(X)\to\mathrm{GL}_{N}(K)$ by
$(\tau_{1}^{ss},\ldots,\tau_{n}^{ss}):\pi_{1}(X)\to\mathrm{GL}_{N_{\tau_{1}}}(K)\times\cdots\times\mathrm{GL}_{N_{\tau_{n}}}(K)\subset\mathrm{GL_{N}(K)}.$
Then $\varrho$ is semisimple. For every connected Zariski closed subset $T$ of $X$, $\sigma(T)$ is a point iff $s_{\tau_{i}}(T)$ is a point for every $i=1,\ldots,n$. This happens iff $\tau_{i}^{ss}({\rm Im}[\pi_{1}(T)\to\pi_{1}(X)])$ is a bounded subgroup of $\mathrm{GL}_{N_{\tau_{i}}}(K)$ for every $i=1,\ldots,n$ (cf. Remark˜1.4). The later is equivalent to the boundedness of $\varrho({\rm Im}[\pi_{1}(T)\to\pi_{1}(X)])$. Hence the map $\sigma:X\to\Sigma$ is a Katzarkov-Eyssidieux reduction map for $\varrho$. Note that $\mathrm{ker}(\varrho)=\mathrm{ker}(\tau_{1}^{ss})\cap\cdots\cap\mathrm{ker}(\tau_{n}^{ss})$. Hence $\cap_{\tau\in\Lambda}\mathrm{ker}(\tau^{ss})\subset\mathrm{ker}(\varrho)$. ∎
1.2 Some facts on algebraic groups
Note that an algebraic group $G$ over a field of characteristic zero is reductive if and only if $G$ is linearly reductive, i.e, for every finite dimensional representation of $G$ is semisimple (cf. [Mil17, Corollary 22.43]). However, this fact fails for algebraic group defined over positive characteristic field. We recall the following example in [Mil17, Example 12.55].
Example 1.7.
Let $k=\overline{\mathbb{F}_{2}}$ and let $V$ be the standard 2-dimensional representation of $\mathrm{SL}_{2}(k)$. Then $\operatorname{Sym}^{2}V$ is not semisimple as a representation. Indeed, let $e_{1}:=(1,0)$ and $e_{2}:=(0,1)$. Within the basis $\{e_{1}e_{2},e_{1}^{2},e_{2}^{2}\}$ of $\operatorname{Sym}^{2}k^{2}$, we can express $\operatorname{Sym}^{2}V$ in the matrix form as
$\mathrm{SL}_{2}(k)$ $\rightarrow\mathrm{GL}_{3}(k)$
$\left(\begin{array}[]{ll}a&b\\ c&d\end{array}\right)$ $\mapsto\left(\begin{array}[]{ccc}1&0&0\\ ab&a^{2}&b^{2}\\ cd&c^{2}&d^{2}\end{array}\right)$
It is easy to see that it is not a semisimple representation since the ${\rm SL}_{2}$-invariant subspace spanned by $\{e_{1}^{2},e_{2}^{2}\}$ have no ${\rm SL}_{2}$-equivariant complement because $ab$ and $cd$ are not linear polynomials in $a^{2},b^{2},c^{2},d^{2}$.
It is worthwhile mentioning that an algebraic group $G$ over a field of characteristic $p\neq 0$ is linearly reductive if and only if its identity component $G^{\circ}$ is a torus and $p$ does not divide the index $\left(G:G^{\circ}\right)$ (cf. [Mil17, Remark 12.56]). Therefore, in this paper, we must distinguish between semisimple and reductive representations as we are working over linear algebraic groups over fields of positive characteristic.
2 Shafarevich morphism in positive characteristic
In this section we will prove Theorem˜A.
2.1 A lemma on finite groups
Lemma 2.1.
Let $K$ be an algebraically closed field of positive characteristic and let $\Gamma$ be a finitely generated group. Let $\varrho:\Gamma\to\mathrm{GL}_{N}(K)$ be a representation such that its semisimplification has finite image. Then $\varrho(\Gamma)$ is finite.
Démonstration.
Since the semisimplification $\varrho^{ss}$ of $\varrho$ has finite image, we can replace $\Gamma$ by a finite index subgroup such that $\varrho^{ss}(\Gamma)$ is trivial. Therefore, some conjugation $\sigma$ of $\varrho$ has image in the subgroup ${\rm U}_{N}(K)$ consisting of all upper-triangular matrices in ${\rm GL}_{N}(K)$ with 1’s on the main diagonal.
Note ${\rm U}_{N}(K)$ admits a central normal series whose successive quotients are isomorphic to $\mathbb{G}_{a,K}$. We remark that a finitely generated subgroup of $\mathbb{G}_{a,K}$ is a finite group, for $K$ has positive characteristic. By [ST00, Proposition 4.17], any finite index subgroup of a finitely generated group is also finitely generated. Consequently, $\sigma(\Gamma)$ admits a central normal series whose successive quotients are finitely generated subgroups of $\mathbb{G}_{a,K}$, which are all finite groups. It follows that $\sigma(\Gamma)$ is finite. The lemma is proved. ∎
2.2 Consideration of character varieties
We first briefly explain the character varieties for finitely generated groups in positive characteristic and refer the readers to [LM85, Ses77] for more details. Let $\Gamma$ be a finitely generated group. The variety of $N$-dimensional linear representations of $\Gamma$ in characteristic zero is represented by an affine $\mathbb{Z}$-scheme $R(\Gamma,N)$ of finite type. Namely, given a commutative ring $A$, the set of $A$-points of $R(\Gamma,N)$ is :
$R(\Gamma,N)(A)={\rm Hom}(\Gamma,{\rm GL}_{N}(A)).$
Let $p$ be a prime number. Consider $R(\Gamma,N)_{\mathbb{F}_{p}}:=R(\Gamma,N)\times_{\textrm{Spec}\,\mathbb{Z}}\textrm{Spec}\,\mathbb{F}_{p}$ and note that the general linear group over $\mathbb{F}_{p}$, denoted by ${\rm GL}(N,\mathbb{F}_{p})$, acts on $R(\Gamma,N)_{\mathbb{F}_{p}}$ by conjugation. Using Seshadri’s extension of geometric invariant theory quotients for schemes of arbitrary field [Ses77, Theorem 3], we can take the GIT quotient of $R(\Gamma,N)_{\mathbb{F}_{p}}$ by ${\rm GL}(N,\mathbb{F}_{p})$, denoted by $M_{\rm B}(\Gamma,N)_{\mathbb{F}_{p}}$. Then $M_{\rm B}(\Gamma,N)_{\mathbb{F}_{p}}$ is also an affine $\mathbb{F}_{p}$-scheme of finite type. For any algebraically closed field $K$ of characteristic $p$, the $K$-points $M_{\rm B}(\Gamma,N)_{\mathbb{F}_{p}}(K)$ is identified with the conjugacy classes of semi-simple representations $\Gamma\to{\rm GL}_{N}(K)$. Namely we have the following.
Proposition 2.2.
Let $K$ be an algebraically closed field of characteristic $p$. Then :
  1. (a)
    Given a linear representation $\varrho:\Gamma\to\mathrm{GL}_{N}(K)$, we have $\varrho^{ss}]=[\varrho]$.
  2. (b)
    For each point $x\in M_{\rm B}(\Gamma,N)_{\mathbb{F}_{p}}(K)$, there exists a semisimple representation
    $\varrho:\Gamma\to\mathrm{GL}_{N}(K)$
    such that $\varrho]=x$.
  3. (c)
    Let $\varrho:\Gamma\to\mathrm{GL}_{N}(K)$ and $\varrho^{\prime}:\Gamma\to\mathrm{GL}_{N}(K)$ be two semisimple representations such that $\varrho]=[\varrho^{\prime}]$. Then $\varrho$ and $\varrho^{\prime}$ are conjugate.
Démonstration.
These are well-known when the field is zero characteristic (cf. [LM85, Thm 1.28]). The proofs also work for positive characteristic case as well, as we briefly present below.
Let $\varrho\in R(\Gamma,N)_{\mathbb{F}_{p}}(K)$ be a linear representation. We denote by $O(\varrho)$ the orbit of $\varrho$ by the conjugacy action, which is a constructible subset of $R(\Gamma,N)\times_{\textrm{Spec}\,\mathbb{Z}}\textrm{Spec}\,K$. Let $\overline{O(\varrho)}$ be the closure of $O(\varrho)$. Then we have $\varrho^{ss}$ is a closed point of $\overline{O(\varrho)}$, whose proof is the same as [LM85, Lem 1.26]. For the GIT quotient $\pi:R(\Gamma,N)_{\mathbb{F}_{p}}\to M_{\rm B}(\Gamma,N)_{\mathbb{F}_{p}}$, we know that $\overline{O(\varrho)}\subset\pi^{-1}([\varrho])$. Therefore, we have $\varrho]=[\varrho^{ss}]$.
Next, let $x\in M_{\rm B}(\Gamma,N)_{\mathbb{F}_{p}}(K)$. Since the GIT quotient is surjective, so is
$\pi_{K}:R(\Gamma,N)_{\mathbb{F}_{p}}(K)\to M_{\rm B}(\Gamma,N)_{\mathbb{F}_{p}}(K).$
Hence we may take $\varrho_{0}\in R(\Gamma,N)_{\mathbb{F}_{p}}(K)$ such that $\varrho_{0}]=x$. Set $\varrho=(\varrho_{0})^{ss}$. Then $\varrho$ is semisimple and $\varrho]=[\varrho_{0}]=x$.
Finally, for any matrix $A\in{\rm GL}_{N}(K)$, we denote by $\chi(A)=T^{N}+\sigma_{1}(A)T^{N-1}+\cdots+\sigma_{N}(A)$ its characteristic polynomial. Given $\gamma\in\Gamma$, we define a map $f_{\gamma}:R(\Gamma,N)_{\mathbb{F}_{p}}(K)\to\mathbb{A}^{N}(K)$ by $f_{\gamma}(\tau)=(\sigma_{1}(\tau(\gamma)),\ldots,\sigma_{N}(\tau(\gamma)))$, where $\tau\in R(\Gamma,N)_{\mathbb{F}_{p}}(K)$. Note that $f_{\gamma}$ is induced by a morphism of $\mathbb{F}_{p}$-schemes, which we still denote by $f_{\gamma}:R(\Gamma,N)_{\mathbb{F}_{p}}\to\mathbb{A}^{N}_{\mathbb{F}_{p}}$. Note that $f_{\gamma}$ is ${\rm GL}(N,\mathbb{F}_{p})$ invariant. Hence $f_{\gamma}$ factors the GIT quotient $\pi:R(\Gamma,N)_{\mathbb{F}_{p}}\to M_{\rm B}(\Gamma,N)_{\mathbb{F}_{p}}$. Now let $\varrho,\varrho^{\prime}\in R(\Gamma,N)_{\mathbb{F}_{p}}(K)$ be two semisimple representations such that $\varrho]=[\varrho^{\prime}]$. Then we have $f_{\gamma}(\varrho)=f_{\gamma}(\varrho^{\prime})$ for all $\gamma\in\Gamma$. In other words, $\varrho(\gamma)$ and $\varrho^{\prime}(\gamma)$ have the same characteristic polynomials for all $\gamma\in\Gamma$. Hence by the Brauer-Nesbitt theorem, the two semisimple representations $\varrho$ and $\varrho^{\prime}$ are conjugate. ∎
Lemma 2.3.
Let $M\subset M_{\rm B}(\Gamma,N)_{\mathbb{F}_{p}}$ be a constructible set. Assume that, for every reductive representation $\tau:\Gamma\to{\rm GL}_{N}(K)$ with $K$ a local field of characteristic $p$ such that $\tau]\in M(K)$, the image $\tau(\Gamma)\subset{\rm GL}_{N}(K)$ is bounded. Then $M$ is zero dimensional.
Démonstration.
Let $\pi:R(\Gamma,N)_{\mathbb{F}_{p}}\to M_{\rm B}(\Gamma,N)_{\mathbb{F}_{p}}$ be the GIT quotient, which is a surjective $\mathbb{F}_{p}$-morphism. Let $T\subset\pi^{-1}(M)$ be any irreducible affine curve defined over $\overline{\mathbb{F}_{p}}$. We shall show that $\pi(T)$ is a point. This proves that $M$ is zero dimensional.
To prove that $\pi(T)$ is a point, we take $\bar{C}$ as the compactification of the normalization $C$ of $T$, and let $\{P_{1},\ldots,P_{\ell}\}=\bar{C}\setminus C$. There exists $q=p^{n}$ for some $n\in\mathbb{Z}_{>0}$ such that $\bar{C}$ is defined over $\mathbb{F}_{q}$ and $P_{i}\in\bar{C}(\mathbb{F}_{q})$ for each $i$. By the universal property of the representation scheme $R(\Gamma,N)$, $C$ gives rise to a representation $\varrho_{C}:\Gamma\to{\rm GL}_{N}(\mathbb{F}_{q}[C])$, where $\mathbb{F}_{q}[C]$ is the coordinate ring of $C$. Consider the discrete valuation $v_{i}:\mathbb{F}_{q}(C)\to\mathbb{Z}$ defined by $P_{i}$, where $\mathbb{F}_{q}(C)$ is the function field of $C$. Let $\widehat{\mathbb{F}_{q}(C)}_{v_{i}}$ be the completion of $F_{q}(C)$ with respect to $v_{i}$. Then we have the isomorphism $\big(\widehat{\mathbb{F}_{q}(C)}_{v_{i}},v_{i}\big)\simeq\big(\mathbb{F}_{q}((t)),v\big)$, where $\big(\mathbb{F}_{q}((t)),v\big)$ is the formal Laurent field of $\mathbb{F}_{q}$ with the valuation $v$ defined by $v(\sum_{i=m}^{+\infty}a_{i}t^{i})=\min\{i\mid a_{i}\neq 0\}$. Let $\varrho_{i}:\Gamma\to{\rm GL}_{N}(\mathbb{F}_{q}((t)))$ be the extension of $\varrho_{C}$ with respect to $\widehat{\mathbb{F}_{q}(C)}_{v_{i}}$. Then we have
$\varrho_{i}]\in M({\mathbb{F}_{q}((t))}).$
Hence, by our assumption, $\varrho_{i}^{ss}(\Gamma)$ is bounded for each $i$. Hence, by Claim˜1.3, $\varrho_{i}(\Gamma)$ is bounded for each $i$. Thus after we replace $\varrho_{i}$ by some conjugation, we have $\varrho_{i}(\Gamma)\subset{\rm GL}_{N}(\mathbb{F}_{q}[[t]])$, where the $\mathbb{F}_{q}[[t]]$ is the ring of integers of $\mathbb{F}_{q}((t))$, i.e.
$\mathbb{F}_{q}[[t]]:=\{\sum_{i=0}^{+\infty}a_{i}t^{i}\mid a_{i}\in\mathbb{F}_{q}\}.$
For any matrix $A\in{\rm GL}_{N}(K)$, we denote by $\chi(A)=T^{N}+\sigma_{1}(A)T^{N-1}+\cdots+\sigma_{N}(A)$ its characteristic polynomial. Since we have assumed that $\varrho_{i}(\Gamma)\subset{\rm GL}_{N}(\mathbb{F}_{q}[[t]])$ for each $i$, it follows that $\sigma_{j}(\varrho_{i}(\gamma))\in\mathbb{F}_{q}[[t]]$ for each $i$. Therefore, by the definition of $\varrho_{i}$, $v_{i}\big(\sigma_{j}(\varrho_{C}(\gamma))\big)\geq 0$ for each $i$. It follows that $\sigma_{j}(\varrho_{C}(\gamma))$ extends to a regular function on $\overline{C}$, which is thus constant. This implies that for any two representations $\eta_{1}:\Gamma\to{\rm GL}_{N}(K_{1})$ and $\eta_{2}:\Gamma\to{\rm GL}_{N}(K_{2})$ such that ${\rm char}\,K_{1}={\rm char}\,K_{2}=p$ and $\eta_{i}\in C(K_{i})$, we have $\chi(\eta_{1}(\gamma))=\chi(\eta_{2}(\gamma))$ for each $\gamma\in\Gamma$. In other words, $\eta_{1}$ and $\eta_{2}$ has the same characteristic polynomial. It follows that $\eta_{1}]=[\eta_{2}]$ by the Brauer-Nesbitt theorem. Hence $\pi(T)$ is a point. Thus $M$ is zero dimensional. ∎
Corollary 2.4.
Let us consider $M\subset M_{\rm B}(X,N)_{\mathbb{F}_{p}}$ satisfying the same assumptions as in Lemma˜2.3. Let $\varrho:\Gamma\to{\rm GL}_{N}(L)$ be a linear representation with $L$ a field of characteristic $p$ and $\varrho]\in M(L)$. Then the image $\varrho(\Gamma)\subset{\rm GL}_{N}(L)$ is finite.
Démonstration.
By Lemma˜2.3, $M$ is zero dimensional. Thus we can find a point $\eta:\Gamma\to{\rm GL}_{N}(\overline{\mathbb{F}_{p}})$ such that $\eta]\in M(\overline{\mathbb{F}_{p}})$ and $\eta]=[\varrho]$ in $M(L)$. Since $\eta(\Gamma)$ is finite, the semisimplification $\eta^{ss}$ of $\eta$ has also finite image. As the semisimplification of $\varrho$ is isomorphic to $\eta^{ss}$, by virtue of Lemma˜2.1, we conclude that $\varrho(\Gamma)$ is finite. ∎
Lemma 2.5.
Let $\varphi:\Gamma^{\prime}\to\Gamma$ be a group morphism from another finitely generated group $\Gamma^{\prime}$. Let $M\subset M_{\rm B}(\Gamma,N)_{\mathbb{F}_{p}}$ be a constructible set. Then the following two statements are equivalent :
  1. (a)
    For every reductive representation $\tau:\Gamma\to{\rm GL}_{N}(K)$ with $K$ a local field of characteristic $p$ such that $\tau]\in M(K)$, the image $\tau\circ\varphi(\Gamma^{\prime})\subset{\rm GL}_{N}(K)$ is bounded.
  2. (b)
    For every linear representation $\varrho:\Gamma\to{\rm GL}_{N}(L)$ with $L$ a field of characteristic $p$ and $\varrho]\in M(L)$, the image $\varrho\circ\varphi(\Gamma^{\prime})\subset{\rm GL}_{N}(L)$ is finite.
Démonstration.
The implication (b) $\implies$ (a) is trivial. In the following we prove the implication (a) $\implies$ (b). We have a induced map $\iota:M_{\rm B}(\Gamma,N)_{\mathbb{F}_{p}}\to M_{\rm B}(\Gamma^{\prime},N)_{\mathbb{F}_{p}}$. Set $M^{\prime}=\iota(M)$. Then $M^{\prime}\subset M_{\rm B}(\Gamma^{\prime},N)_{\mathbb{F}_{p}}$ is a constructible set. We shall show that $M^{\prime}$ satisfies the assumption of Lemma˜2.3. Indeed let $\sigma:\Gamma\to{\rm GL}_{N}(K^{\prime})$ be a reductive representation such that $\sigma]\in M^{\prime}(K^{\prime})$, where $K^{\prime}$ is a local field of characteristic $p$. Then we may take a reductive representation $\tau:\Gamma\to{\rm GL}_{N}(\overline{K^{\prime}})$ such that $\tau]\in M(\overline{K^{\prime}})$ and $\iota([\tau])=[\sigma]$. By the assumption (a), $\tau\circ\varphi(\Gamma^{\prime})$ is bounded. By $\iota([\tau])=[\tau\circ\varphi]$, we have $\tau\circ\varphi]=[\sigma]$. Thus by Claim˜1.3, $\sigma(\Gamma^{\prime})$ is bounded. Thus $M^{\prime}$ satisfies the assumption of Lemma˜2.3.
Now let $\varrho:\Gamma\to{\rm GL}_{N}(L)$ be a linear representation such that $\varrho]\in M(L)$ with $L$ a field of characteristic $p$. Then we have $\varrho\circ\varphi]\in M^{\prime}(L)$. By Corollary˜2.4, $\varrho\circ\varphi(\Gamma^{\prime})$ is finite. ∎
2.3 Factorization through non-rigidity
Let $X$ be a quasi-projective smooth variety. We write $M_{\rm B}(X,N)_{\mathbb{F}_{p}}=M_{\rm B}(\pi_{1}(X),N)_{\mathbb{F}_{p}}$. Let $M\subset M_{\rm B}(X,N)_{\mathbb{F}_{p}}$ be a Zariski closed subset.
Definition 2.6.
The reduction map $s_{M}:X\to S_{M}$ is obtained through the simultaneous Stein factorization of the reductions $\{s_{\tau}:X\to S_{\tau}\}_{[\tau]\in M(K)}$. Here $\tau:\pi_{1}(X)\to{\rm GL}_{N}(K)$ ranges over all reductive representations with $K$ a non-archimedean local field of characteristic $p$ such that $\tau]\in M(K)$ and $s_{\tau}:X\to S_{\tau}$ is the reduction map defined in Theorem˜1.2.
The reduction map $s_{M}:X\to S_{M}$ enjoys the following crucial property.
Theorem 2.7.
Let $M$ be a Zariski closed subset of $M_{\rm B}(X,N)_{\mathbb{F}_{p}}$. The reduction map $s_{M}:X\to S_{M}$ is the Shafarevich morphism for $M$. That is, for any connected Zariski closed subset $Z$ of $X$, the following properties are equivalent :
  1. (a)
    $s_{M}(Z)$ is a point ;
  2. (b)
    for any linear representation $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ with $K$ a field of characteristic $p$ and $\varrho]\in M(K)$, we have $\varrho({\rm Im}[\pi_{1}(Z)\to\pi_{1}(X)])$ is finite ;
  3. (c)
    for any semisimple representation $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ with $K$ a field of characteristic $p$ such that $\varrho]\in M(K)$, we have $\varrho({\rm Im}[\pi_{1}(Z_{o}^{\rm norm})\to\pi_{1}(X)])$ is finite, where $Z_{o}$ is any irreducible component of $Z$.
Démonstration.
(a) $\implies$ (b) : Let $\tau:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be a reductive representation with $K$ a local field of characteristic $p$ such that $\tau]\in M(K)$. Then by the assumption (a), $s_{\tau}(Z)$ is a point. Hence $\tau({\rm Im}[\pi_{1}(Z)\to\pi_{1}(X)])$ is bounded. Hence by Lemma˜2.5, $\varrho({\rm Im}[\pi_{1}(Z)\to\pi_{1}(X)])$ is finite for every linear representation $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ with $K$ a field of characteristic $p$ and $\varrho]\in M(K)$.
(b) $\implies$ (c) : this is obvious.
(c) $\implies$ (a) : Let $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ with $K$ any non-archimedean local field of characteristic $p$ such that $\varrho]\in M(K)$. Then, the image $\varrho({\rm Im}[\pi_{1}(Z_{o}^{\rm norm})\to\pi_{1}(X)])$ is finite by our assumption, and is thus bounded. By the property in Theorem˜1.2, $s_{\varrho}(Z)$ is a point. By Definition˜2.6, $s_{M}(Z)$ is also a point. ∎
Remark 2.8.
We remark that in the case of characteristic 0, i.e. meaning that $M$ is a Zariski closed subset of $M_{\rm B}(X,N)$ defined over $\bar{\mathbb{Q}}$, the reduction map $s_{M}:X\to S_{M}$ constructed in [DYK23] is the same as Definition˜2.6 : it is obtained through the simultaneous Stein factorization of the reductions $\{s_{\tau}:X\to S_{\tau}\}_{[\tau]\in M(K)}$, where $\tau:\pi_{1}(X)\to{\rm GL}_{N}(K)$ ranges over all reductive representations with $K$ a local field of characteristic 0 such that $\tau]\in M(K)$ and $s_{\tau}:X\to S_{\tau}$ is the reduction map defined in Theorem˜1.2. The interested readers can refer to [DYK23, §3.1] for further details.
2.4 Construction of the Shafarevich morphism
Theorem 2.9.
Let $X$ be a quasi-projective normal variety and $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be a linear representation, where $K$ is a field of characteristic $p>0$. Then the Shafarevich morphism ${\rm sh}_{\varrho}:X\to{\rm Sh}_{\varrho}(X)$ exists. That is, for any connected Zariski closed subset $Z\subset X$, the following properties are equivalent :
  1. (a)
    ${\rm sh}_{\varrho}(Z)$ is a point ;
  2. (b)
    $\varrho({\rm Im}[\pi_{1}(Z)\to\pi_{1}(X)])$ is finite ;
  3. (c)
    for each irreducible component $Z_{o}$ of $Z$, $\varrho^{ss}({\rm Im}[\pi_{1}(Z_{o}^{\rm norm})\to\pi_{1}(X)])$ is finite.
Démonstration.
Step 1 : The smooth case. In this step, we will assume that $X$ is smooth. Define
$M:=\bigcap_{f:Y\to X}j_{f}^{-1}\{[1]\},$ (2.1)
where $1$ stands for the trivial representation, and $f:Y\to X$ ranges over all proper morphisms from positive dimensional quasi-projective normal varieties such that $f^{*}\varrho=1$. Here $j_{f}:M_{\rm B}(X,N)_{\mathbb{F}_{p}}\to M_{\rm B}(Y,N)_{\mathbb{F}_{p}}$ is a morphism of affine $\mathbb{F}_{p}$-scheme induced by $f$. Then $M$ is a Zariski closed subset. We apply Theorem˜2.7 to construct the Shafarevich morphism $s_{M}:X\to S_{M}$ associated with $M$. It is a dominant morphism with general fibers connected. Let ${\rm sh}_{\varrho}:X\to{\rm Sh}_{\varrho}(X)$ be $s_{M}:X\to S_{M}$ and we will prove that it satisfies the properties in the theorem.
(a) $\Rightarrow$ (b) : this follows from the fact that $\varrho]\in M(K)$ and Theorem˜2.7.
(b) $\Rightarrow$ (c) : obvious.
(c) $\Rightarrow$ (a) : We take a finite étale cover $Y\to Z^{\rm norm}_{o}$ such that $f^{*}\varrho^{ss}(\pi_{1}(Y))$ is trivial, where we denote by $f:Y\to X$ the natural proper morphism. Let $\tau:\pi_{1}(X)\to{\rm GL}_{N}(L)$ be any linear representation such that $\tau]\in M(L)$ where $L$ is any field of characteristic $p$. Then $f^{*}\tau]=[1]$ by (2.1). Thanks to Lemma˜2.1, $f^{*}\tau(\pi_{1}(Y))$ is finite, and it follows that $\tau({\rm Im}[\pi_{1}(Z^{\rm norm}_{o})\to\pi_{1}(X)])$ is finite as ${\rm Im}[\pi_{1}(Y)\to\pi_{1}(Z^{\rm norm}_{o})]$ is a finite index subgroup of $\pi_{1}(Z^{\rm norm}_{o})$. According to Theorem˜2.7, $s_{M}(Z)$, and thus ${\rm sh}_{\varrho}(Z)$ is a point.
Step 2 : Reduction to the smooth case. We now do not assume that $X$ is smooth. Let $\mu:X_{1}\to X$ be a resolution of singularities. Then the Shafarevich morphism ${\rm sh}_{\mu^{*}\varrho}:X_{1}\to{\rm Sh}_{\mu^{*}\varrho}(X_{1})$ exists and satisfies the properties in the theorem. Since $X$ is normal, each fiber $F$ of $\mu$ is compact and connected. Since $\mu^{*}\varrho({\rm Im}[\pi_{1}(F)\to\pi_{1}(X_{1})])=\{1\}$, it implies that ${\rm sh}_{\mu^{*}\varrho}(F)$ is a point. Hence there exists a morphism ${\rm sh}_{\varrho}:X\to{\rm Sh}_{\mu^{*}\varrho}(X_{1})$ such that ${\rm sh}_{\varrho}\circ\mu={\rm sh}_{\mu^{*}\varrho}$.
(a) $\Rightarrow$ (b) : Let $W:=\mu^{-1}(Z)$, which is a connected Zariski closed subset of $X_{1}$. Then ${\rm sh}_{\mu^{*}\varrho}(W)$ is a point, which implies that $\mu^{*}\varrho({\rm Im}[\pi_{1}(W)\to\pi_{1}(X_{1})])$ is finite. By [DYK23, Lemma 3.47], $\pi_{1}(W)\to\pi_{1}(Z)$ is surjective. It follows that
$\varrho({\rm Im}[\pi_{1}(Z)\to\pi_{1}(X)])=\mu^{*}\varrho({\rm Im}[\pi_{1}(W)\to\pi_{1}(X_{1})])$
is finite.
(b) $\Rightarrow$ (c) : obvious.
(c) $\Rightarrow$ (a) : Let $W$ be an irreducible component of $\mu^{-1}(Z_{o})$ which is surjective onto $Z_{o}$. This implies that $\mu^{*}\varrho^{ss}({\rm Im}[\pi_{1}(W^{\rm norm})\to\pi_{1}(X_{1})])$ is finite. Since $(\mu^{*}\varrho)^{ss}]=[\mu^{*}\varrho^{ss}]$, by Lemma˜2.1, $(\mu^{*}\varrho)^{ss}({\rm Im}[\pi_{1}(W^{\rm norm})\to\pi_{1}(X_{1})])$ is also finite. Hence ${\rm sh}_{\varrho}(Z_{o})={\rm sh}_{\mu^{*}\varrho}(W)$ is a point. Since $Z$ is connected, we conclude that ${\rm sh}_{\varrho}(Z)$ is a point.
Letting ${\rm Sh}_{\varrho}(X):={\rm Sh}_{\mu^{*}\varrho}(X_{1})$, we prove the theorem. ∎
Lemmas˜1.6, 2.7 and 2.9 yield the following result.
Corollary 2.10.
Let $X$ be a smooth quasi-projective variety and $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be a linear representation, where $K$ is a field of characteristic $p>0$. Then there exist
  • a local field $K^{\prime}$ of $\mathrm{char}\ K^{\prime}=p$,
  • a positive integer $N^{\prime}>0$,
  • a reductive representation $\sigma:\pi_{1}(X)\to\mathrm{GL}_{N^{\prime}}(K^{\prime})$
such that the Katzarkov-Eyssidieux reduction map $s_{\sigma}:X\to S_{\sigma}$ of $\sigma$ is the Shafarevich morphism ${\rm sh}_{\varrho}:X\to{\rm Sh}_{\varrho}(X)$ of $\varrho$.
Démonstration.
By Theorems˜2.7 and 2.9, ${\rm sh}_{\varrho}:X\to{\rm Sh}_{\varrho}(X)$ is obtained through the simultaneous Stein factorization of the reduction maps $\{s_{\tau}:X\to S_{\tau}\}_{[\tau]\in M(K)}$, where
$\tau:\pi_{1}(X)\to{\rm GL}_{N}(K)$
ranges over all reductive representations with $K$ a local field of characteristic $p$, and $M$ is defined in (2.1). By Lemma˜1.6, there exist a local field $K$ of $\mathrm{char}\ K=p$, a positive integer $N^{\prime}$ and a reductive representation $\sigma:\pi_{1}(X)\to\mathrm{GL}_{N^{\prime}}(K)$ such that ${\rm sh}_{\varrho}:X\to{\rm Sh}_{\varrho}(X)$ coincides with $s_{\sigma}:X\to S_{\sigma}$. ∎
3 Hyperbolicity via linear representation in positive characteristic
The structure of the Shafarevich morphism, as presented in the proof of Theorem˜2.9, is related to the hyperbolicity of algebraic varieties. We will prove Theorems˜C, D and E in this section.
3.1 On the generalized Green-Griffiths-Lang conjecture
Theorem 3.1.
Let $X$ be a quasi-projective normal variety. Let $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be a big representation where $K$ is a field of positive characteristic. Then the following properties are equivalent :
  1. (i)
    $X$ is of log general type ;
  2. (ii)
    $X$ is pseudo Picard hyperbolic ;
  3. (iii)
    $X$ is pseudo Brody hyperbolic ;
  4. (iv)
    $X$ is strongly of log general type.
It’s worth noting that the conjectural four equivalent properties mentioned in Theorem˜3.1 constitute the statement of the generalized Green-Griffiths-Lang conjecture, as presented in [CDY22].
Proof of Theorem˜3.1.
[…]
  1. (a)
    there exists a set of forms $\{\eta_{j}\}_{j=1,\ldots,m}\subset H^{0}(\overline{X^{\mathrm{sp}}},\pi^{*}\Omega_{\overline{X}}(\log D))$ which are invariant under $H$ ;
  2. (b)
    $\pi$ is étale outside
    $R:=\{x\in\overline{X^{\mathrm{sp}}}\mid\exists\eta_{j}\neq\eta_{\ell}\mbox{ with }(\eta_{j}-\eta_{\ell})(x)=0\}.$ (3.1)
  3. (c)
    There exists a morphism $a:X^{\mathrm{sp}}\to A$ to a semi-abelian variety $A$ with $H$ acting on $A$ such that $a$ is $H$-equivariant, where $X^{\!\rm sp}:=\pi^{-1}(X)$. Here, $X^{\!\rm sp}\to X$ is called the spectral cover of $X$ associated with $\tau$.
  4. (d)
    The reduction map $s_{\tau}:X\to S_{\tau}$ is the quasi-Stein factorization of the quotient $X\to A/H$ of $a$ by $H$.
Claim 3.2.
We have $\dim X^{\!\rm sp}=\dim a(X^{\!\rm sp})$.
Démonstration.
Since $\varrho$ is big, it follows that ${\rm sh}_{\varrho}$, hence $s_{\tau}$ is birational. Thus $\dim X^{\!\rm sp}=\dim a(X^{\!\rm sp})$ by Item˜(d). ∎
Now we will apply the techniques and results in [CDY22] to prove the theorem.
(i) $\Rightarrow$ (ii) : We first recall some notions in Nevanlinna theory used in [CDY22, §4.1]. Let $Y$ be a connected Riemann surface with a proper surjective holomorphic map $p:Y\to\mathbb{C}_{>\delta}$, where $\mathbb{C}_{>\delta}:=\{z\in\mathbb{C}\mid\delta<|z|\}$ with some fixed positive constant $\delta>0$. For $r>2\delta$, define $Y(r)=p^{-1}\big(\mathbb{C}_{>2\delta}(r)\big)$ where $\mathbb{C}_{>2\delta}(r)=\{z\in\mathbb{C}\mid 2\delta<|z|<r\}$. In the following, we assume that $r>2\delta$. The ramification counting function of the covering $p:Y\to\mathbb{C}_{>\delta}$ is defined by
$N_{{\rm ram}\,p}(r):=\frac{1}{{\rm deg}p}\int_{2\delta}^{r}\left[\sum_{y\in{Y}(t)}{\rm ord}\,_{y}{\rm ram}\,p\right]\frac{dt}{t},$ (3.2)
where ${\rm ram}\,p\subset Y$ is the ramification divisor of $p:Y\to\mathbb{C}_{>\delta}$.
For any holomorphic map $f:\mathbb{C}_{>\delta}\to X$ whose image is not contained in $\pi(R)$, there exists a surjective finite holomorphic map $p:Y\to\mathbb{C}_{>\delta}$ from a connected Riemann surface $Y$ to $\mathbb{C}_{>\delta}$ and a holomorphic map $g:Y\to X^{\!\rm sp}$ satisfying the following diagram :
[…]
Here $Y$ is a connected component of the normalization of $\mathbb{C}_{>\delta}\times_{X}X^{\!\rm sp}$. By [CDY22, Proposition 6.9], there exists a proper Zariski closed subset $E\subsetneq X$ such that for any holomorphic map $f:\mathbb{C}_{>\delta}\to X$ whose image is not contained in $E$, one has
$N_{{\rm ram}\,p}(r)=o(T_{g}(r,L))+O(\log r)$ (3.4)
for all $r>2\delta$ outside some exceptional interval with finite Lebesgue measure. Here $g:Y\to X^{\!\rm sp}$ is the induced holomorphic map in (3.3), $L$ is an ample line bundle on $\overline{X^{\!\rm sp}}$ equipped with a smooth hermitian metric $h_{L}$ and $T_{g}(r,L)$ is the order function defined by
$T_{g}(r,L):=\frac{1}{\operatorname{deg}p}\int_{2\delta}^{r}\left[\int_{Y(t)}g^{*}c_{1}(L,h_{L})\right]\frac{dt}{t}.$ (3.5)
Note that $X^{\!\rm sp}$ is of log general type as we assume that $X$ is of log general type and $\pi:X^{\!\rm sp}\to X$ is a Galois cover. We apply [CDY22, Theorem 4.1] to conclude that there exists a proper Zariski closed set $\Xi\subsetneqq X^{\!\rm sp}$ such that an extension $\overline{g}:\overline{Y}\to\overline{X^{\!\rm sp}}$ of $g$ exists provided $g(Y)\not\subset\Xi$, where $\overline{Y}$ is a Riemann surface such that $p:Y\to\mathbb{C}_{>\delta}$ extends to a proper map $\overline{p}:\overline{Y}\to\mathbb{C}_{>\delta}\cup\{\infty\}$. This induces an extension $\bar{f}:\mathbb{C}_{>\delta}\cup\{\infty\}\to\overline{X}$. Indeed, given a sequence $(a_{n})$ in $\mathbb{C}_{>\delta}$ satisfing $a_{n}\to\infty$, we may take a sequence $(b_{n})$ in $Y$ such that $p(b_{n})=a_{n}$ and $b_{n}\to\beta\in\overline{Y}$. Then $\bar{p}(\beta)=\infty$ and $f(a_{n})=\pi\circ g(b_{n})\to\pi\circ\bar{g}(\beta)$. Thus $(f(a_{n}))$ converges for every sequence $(a_{n})$ satisfing $a_{n}\to\infty$. Thus $f$ has continuous extension $\bar{f}:\mathbb{C}_{>\delta}\cup\{\infty\}\to\overline{X}$, which is holomorphic by Riemann’s removable singularity theorem. Hence, $X$ is pseudo Picard hyperbolic.
(ii) $\Rightarrow$ (iii) : It is obvious that pseudo Picard hyperbolicity implies pseudo Brody hyperbolicity (cf. [CDY22, Lemma 4.3]).
(iii) $\Rightarrow$ (iv) : Since $\pi:X^{\!\rm sp}\to X$ is a finite surjective morphism, $X^{\!\rm sp}$ is also pseudo Brody hyperbolic. By [CDY22, Corollary 4.2], we conclude that there exists a proper Zariski closed subset $Z\subsetneq X^{\!\rm sp}$ such that any positive dimensional closed subvariety $V$ is of log general type if it is not contained in $Z$. Then the rest of the proof is basically the same as that of [CDY22, Theorem 6.3] and let us explain it for the sake of completeness.
By the proof of [CDY22, Theorem 6.3], there exists a Zariski closed subset $\Xi\subset X$ such that we have the following properties :
  1. (1)
    $\pi(Z\cup R)\subset\Xi$, where $R$ is defined in eq.˜3.1 ;
  2. (2)
    Let $V\subset X$ be any closed subvariety such that $V\not\subset\Xi$. Let $W\to V$ be a smooth modification and let $\overline{W}$ be a smooth projective compactification of $W$ such that $D_{\overline{W}}:=\overline{W}-W$ is a simple normal crossing divisor, and $W\to X$ extends to a morphism $\overline{W}\to\overline{X}$. Let $\overline{S}$ be the normalization of an irreducible component of $\overline{W}\times_{\overline{X}}\overline{X^{\!\rm sp}}$.
    $\overline{S}$$\overline{X^{\!\rm sp}}$$\overline{W}$$\overline{X}$$\scriptstyle{g}$$\scriptstyle{p}$$\scriptstyle{\pi}$
    Then by [CDY22, Claim 6.5] the finite morphism $p:\overline{S}\to\overline{W}$ is a Galois morphism with Galois group $H^{\prime}\subset H$. The morphism $g:\overline{S}\to\overline{X^{\!\rm sp}}$ is $H^{\prime}$-equivariant.
  3. (3)
    Define $\psi_{j}:=g^{*}\eta_{j}\in H^{0}(\overline{S},p^{*}\Omega_{\overline{W}}(\log D_{\overline{W}}))$ for $j=1,\ldots,m$. Let $I$ be the set of all $(i,j)$ such that
    • $\eta_{i}-\eta_{j}\neq 0$ ;
    • the image of $S\to X^{\!\rm sp}$ intersects with $\{z\in\overline{X^{\!\rm sp}}\mid(\eta_{i}-\eta_{j})(z)=0\}$.
    Then by by [CDY22, Claim 6.6], for $(i,j)\in I$, $\psi_{i}-\psi_{j}\not=0$ in $H^{0}(\overline{S},p^{*}\Omega_{\overline{W}}(\log D_{\overline{W}}))$.
We set
$R^{\prime}:=\{z\in\overline{S}\mid\exists(i,j)\in I\mbox{ with }(\psi_{i}-\psi_{j})(z)=0\}.$
Then $R^{\prime}$ is a proper Zariski closed subset of $\overline{S}$. Denote by $R_{0}$ the ramification locus of $p:\overline{S}\to\overline{W}$. By the purity of branch locus of finite morphisms, we know that $R_{0}$ is a (Weil) divisor, and thus $p(R_{0})$ is also a divisor. Moreover $R_{0}=p^{-1}\big(p(R_{0})\big)$ since $p$ is Galois with Galois group $H^{\prime}$. Denote by $E$ the sum of prime components of $p(R_{0})$ which intersect $W$. One observes that $S-p^{-1}(E)\to W-E$ is finite étale.
Since $\pi:\overline{X^{\!\rm sp}}\to\overline{X}$ is étale over $\overline{X^{\!\rm sp}}-R$, it follows that $p$ is étale over $\overline{S}-g^{-1}(R)$ and thus $R_{0}\cap S\subset R^{\prime}\cap S.$
Since $\overline{S}\to\overline{X^{\!\rm sp}}$ is $H^{\prime}$-equivariant, it follows that any $h\in H^{\prime}$ acts on
$\{\psi_{i}-\psi_{j}\}_{(i,j)\in I}\subset H^{0}(\overline{S},p^{*}\Omega_{\overline{W}}(\log D_{\overline{W}}))$
as a permutation by our choice of $I$. Define a section
$\sigma:=\prod_{h\in H^{\prime}}\prod_{(i,j)\in I}h^{*}(\psi_{i}-\psi_{j})\in H^{0}(\overline{S},{\text{Sym}}^{\ell}p^{*}\Omega_{\overline{W}}(\log D_{\overline{W}})),$
which is non-zero and vanishes at $R^{\prime}$ by our choice of $I$. Then it is invariant under the $H^{\prime}$-action and thus descends to a section
$\sigma^{H^{\prime}}\in H^{0}(\overline{W},{\text{Sym}}^{\ell}\Omega_{\overline{W}}(\log D_{\overline{W}}))$
so that $p^{*}\sigma^{H^{\prime}}=\sigma$. Since $R_{0}\cap S\subset R^{\prime}\cap S$ and $p^{-1}(p(R_{0}))=R_{0}$, $\sigma^{H^{\prime}}$ vanishes at the divisor $E$. This implies that there is a non-trivial morphism
$\mathcal{O}_{\overline{W}}(E)\to{\text{Sym}}^{\ell}\Omega_{\overline{W}}(\log D_{\overline{W}}).$ (3.6)
[…]
(iv) $\Rightarrow$ (i) : This is obvious. Thus we have completed the proof of Theorem˜3.1. ∎
Remark 3.3.
From the proof of (i) $\Rightarrow$ (ii) in Theorem˜3.1, we note that the weaker condition of $X^{\!\rm sp}$ being of log general type is sufficient to establish the pseudo Picard hyperbolicity of $X$. This observation will be particularly relevant for the proofs of Proposition˜3.5 and Theorem˜3.7 below.
3.2 Comparison of special subsets
This subsection is devoted to the proof of Theorem˜D. We first prove a lemma on the special subset of the big representation in positive characteristic.
Lemma 3.4.
Let $X$ be a smooth quasi-projective variety. Let $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be a big representation where $K$ is a field of positive characteristic. Then the special subset ${\rm Sp}(\varrho)$ defined in Definition˜0.1 is a proper Zariski closed subset of $X$.
Démonstration.
Note that the Shafarevich morphism ${\rm sh}_{\varrho}:X\to{\rm Sh}_{\varrho}(X)$ exists, which is dominant morphism with connected general fibers (cf. Theorem˜2.9). Since $\varrho$ is big, it follows that ${\rm sh}_{\varrho}$ is birational. Therefore, we have a Zariski dense open set $X^{\circ}\subset X$ such that ${\rm sh}_{\varrho}:X^{\circ}\to{\rm sh}_{\varrho}(X^{\circ})$ is an isomorphism. By Theorem˜2.9, for any positive dimensional closed subvariety $Z\subset X$ not contained in $X\setminus X^{\circ}$, $\varrho({\rm Im}[\pi_{1}(Z)\to\pi_{1}(X)])$ is infinite. Hence ${\rm Sp}(\varrho)\subset X\setminus X^{\circ}$. This concludes that ${\rm Sp}(\varrho)$ is a proper Zariski closed subset of $X$. ∎
Before proceeding to the proof of Theorem˜D (see Theorem˜3.6 below), we will first prove the following result.
Proposition 3.5.
Let $X$ be a smooth quasi-projective variety. Let $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be a big representation where $K$ is a field of positive characteristic. Assume that there exists a holomorphic map $f:\mathbb{D}^{*}\to X$ such that $f$ has essential singularity at the origin and that the image $f(\mathbb{D}^{*})$ is Zariski dense. Then $\mathrm{Sp}_{\mathrm{sab}}(X)=X$.
Démonstration.
We will maintain the same notations as introduced in the proof of Theorem˜3.1. Let $\pi:X^{\!\rm sp}\to X$ be the Galois covering, as defined therein. By the assumption, $X$ is not pseudo Picard hyperbolic. Hence $X^{\!\rm sp}$ is not of log general type by Remark˜3.3. Recall that in the proof of Theorem˜3.1, we have a morphism $a:X^{\!\rm sp}\to A$, where $A$ is a semiabelian variety, and $\dim X^{\!\rm sp}=\dim a(X^{\!\rm sp})$. Note that $\bar{\kappa}(X^{\!\rm sp})\geq 0$. We denote by $j:X^{\!\rm sp}\dashrightarrow J(X^{\!\rm sp})$ the logarithmic Iitaka fibration of $X^{\!\rm sp}$, whose general fibers are positive dimensional as $X^{\!\rm sp}$ is not of log general type. After replacing $X^{\!\rm sp}$ be a proper birational model, we might assume that $X^{\!\rm sp}$ is smooth and $j$ is regular. Then for a very general fiber $F$ of $j$, we have $\bar{\kappa}(F)=0$ and $\dim F=\dim a(F)>0$. By [CDY22, Lemma 3.5], we have $\mathrm{Sp}_{\mathrm{sab}}(F)=F$. Consequently, $\mathrm{Sp}_{\mathrm{sab}}(X^{\!\rm sp})=X^{\!\rm sp}$. Since $\pi:X^{\!\rm sp}\to X$ is surjective, it follows that $\mathrm{Sp}_{\mathrm{sab}}(X)=X$. The proposition is proved. ∎
Theorem 3.6 (=Theorem˜D).
Let $X$ be a quasi-projective normal variety. Let
$\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$
be a big representation where $K$ is a field of positive characteristic. Then we have
$\mathrm{Sp}_{\mathrm{sab}}(X)\setminus{\rm Sp}(\varrho)=\mathrm{Sp}_{\mathrm{alg}}(X)\setminus{\rm Sp}(\varrho)=\mathrm{Sp}_{\mathrm{p}}(X)\setminus{\rm Sp}(\varrho)=\mathrm{Sp}_{\mathrm{h}}(X)\setminus{\rm Sp}(\varrho).$
We have ${\rm Sp}_{\bullet}(X)\subsetneq X$ if and only if $X$ is of log general type, where $\rm Sp$$_{\bullet}$ denotes any of $\mathrm{Sp}_{\mathrm{sab}}$, $\mathrm{Sp}_{\mathrm{alg}}$, $\mathrm{Sp}_{\mathrm{h}}$ or $\mathrm{Sp}_{\mathrm{p}}$.
Démonstration.
By [CDY22, Lemma 4.3] one has
$\mathrm{Sp}_{\mathrm{sab}}(X)\subseteq\mathrm{Sp}_{\mathrm{h}}(X)\subseteq\mathrm{Sp}_{\mathrm{p}}(X).$ (3.7)
Let $f:\mathbb{D}^{*}\to X$ be a holomorphic map with essential singularity at the origin such that $f(\mathbb{D}^{*})\not\subset{\rm Sp}(\varrho)$. Let $Z$ be a desingularization of the Zariski closure of $f(\mathbb{D}^{*})$. By the definition of ${\rm Sp}(\varrho)$ in Definition˜0.1, we note that the natural morphism $\iota:Z\to X$ induces a big representation $\iota^{*}\varrho$. Since $\mathrm{Sp}_{\mathrm{p}}(Z)=Z$, Theorem˜3.1 implies that $Z$ is not of log general type. Hence $\iota(Z)\subset\mathrm{Sp}_{\mathrm{alg}}(X)$, which implies
$\mathrm{Sp}_{\mathrm{p}}(X)\setminus{\rm Sp}(\varrho)\subseteq\mathrm{Sp}_{\mathrm{alg}}(X)\setminus{\rm Sp}(\varrho).$ (3.8)
By Proposition˜3.5, $\mathrm{Sp}_{\mathrm{sab}}(Z)=Z$. It follows that
$\mathrm{Sp}_{\mathrm{p}}(X)\setminus{\rm Sp}(\varrho)\subseteq\mathrm{Sp}_{\mathrm{sab}}(X)\setminus{\rm Sp}(\varrho).$
Combining this with (3.7), we get
$\mathrm{Sp}_{\mathrm{sab}}(X)\setminus{\rm Sp}(\varrho)=\mathrm{Sp}_{\mathrm{h}}(X)\setminus{\rm Sp}(\varrho)=\mathrm{Sp}_{\mathrm{p}}(X)\setminus{\rm Sp}(\varrho).$
Thus it remains to prove $\mathrm{Sp}_{\mathrm{p}}(X)\setminus{\rm Sp}(\varrho)=\mathrm{Sp}_{\mathrm{alg}}(X)\setminus{\rm Sp}(\varrho)$.
Let $Y$ be a closed subvariety of $X$ that is not of log general type. Assume that $Y\not\subseteq{\rm Sp}(\varrho)$. Let $\iota:Z\to Y$ be a desingularization. Then $\iota^{*}\varrho$ is a big representation. By Theorem˜3.1, we have $\mathrm{Sp}_{\mathrm{p}}(Y)=Y$. Hence $Y\subset\mathrm{Sp}_{\mathrm{p}}(X)$, which implies
$\mathrm{Sp}_{\mathrm{alg}}(X)\setminus{\rm Sp}(\varrho)\subseteq\mathrm{Sp}_{\mathrm{p}}(X)\setminus{\rm Sp}(\varrho).$
Together with (3.8), we obtain $\mathrm{Sp}_{\mathrm{p}}(X)\setminus{\rm Sp}(\varrho)=\mathrm{Sp}_{\mathrm{alg}}(X)\setminus{\rm Sp}(\varrho)$. We have proved our theorem. ∎
3.3 A characterization of hyperbolicity via fundamental groups
Theorem 3.7.
Let $X$ be a quasi-projective normal variety. Let $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be a big representation where $K$ is a field of positive characteristic. If the Zariski closure $G$ of $\varrho(\pi_{1}(X))$ is a semisimple algebraic group, then ${\rm Sp}_{\bullet}(X)\subsetneq X$ where $\rm Sp$$_{\bullet}$ denotes any of $\mathrm{Sp}_{\mathrm{sab}}$, $\mathrm{Sp}_{\mathrm{alg}}$, $\mathrm{Sp}_{\mathrm{h}}$ or $\mathrm{Sp}_{\mathrm{p}}$, in particular, $X$ is of log general type.
Démonstration.
We may assume that $K$ is algebraically closed. Replacing $X$ by a desingularization, we may assume that $X$ is smooth. We will still maintain the same notations as introduced in the proof of Theorem˜3.1. Let $\pi:X^{\!\rm sp}\to X$ be the Galois covering defined therein. Consider the representation $\pi^{*}\varrho:\pi_{1}(X^{\!\rm sp})\to G(K)$, which is Zariski dense as ${\rm Im}[\pi_{1}(X^{\!\rm sp})\to\pi_{1}(X)]$ is a finite index subgroup of $\pi_{1}(X)$. By the proof of Theorem˜3.1, there exists a morphism $a:X^{\!\rm sp}\to A$ where $A$ is a semiabelian variety such that $\dim X^{\!\rm sp}=\dim a(X^{\!\rm sp})$. Hence we have $\bar{\kappa}(X^{\!\rm sp})\geq 0$.
Claim 3.8.
$X^{\!\rm sp}$ is of log general type.
Démonstration.
Let $\mu:Y\to X^{\!\rm sp}$ be a desingularization such that the logarithmic Iitaka fibration $j:Y\to J(Y)$ is regular. For a very general fiber $F$ of $j$, we have $\bar{\kappa}(F)=0$ and $\dim F=\dim a(F)$. By [CDY22, Lemma 3.3], we have $\pi_{1}(F)$ is abelian.
We write $\tau=(\pi\circ\mu)^{*}\varrho:\pi_{1}(Y)\to G(K)$. Notably, $\tau(\pi_{1}(Y))$ is Zariski dense in $G$. By [Kol93, Corollary 2.9.3] (or [CDY22, Lemma 2.2]), ${\rm Im}[\pi_{1}(F)\to\pi_{1}(Y)]$ is a normal subgroup of $\pi_{1}(Y)$. Consequently, the Zariski closure $N$ of $\tau({\rm Im}[\pi_{1}(F)\to\pi_{1}(Y)])$ is a normal subgroup of $G$. It’s worth noting that the connected component $N^{\circ}$ of $N$ is a tori since $\tau({\rm Im}[\pi_{1}(F)\to\pi_{1}(Y)])$ is commutative. Therefore, $N^{\circ}$ must be trivial since $G$ is assumed to be semisimple. Consequently, $\tau({\rm Im}[\pi_{1}(F)\to\pi_{1}(Y)])$ is finite.
Given our assumption that $\varrho$ is big, we conclude that $\tau$ is also big. Therefore, we arrive at the conclusion that $\dim F=0$, leading us to deduce that both $Y$ and, consequently, $X^{\!\rm sp}$ are of log general type. ∎
[…]
Remark 3.9.
Note that the condition in Theorem˜3.7 is sharp. For example, the representation
$\mathbb{Z}\simeq\pi_{1}(\mathbb{C}^{*})$ $\to{\rm GL}_{1}(\mathbb{F}_{p}(t))$
$n$ $\mapsto t^{n}$
is a big and Zariski dense representation. However, $\mathbb{C}^{*}$ is not of general type and contains a Zariski dense entire curve. This example demonstrates that the semisimplicity of $G$ is necessary for Theorem˜3.7 to hold.
Furthermore, the condition of bigness in Theorem˜3.7 is also indispensable. Indeed, it’s worth noting that for any complex algebraic variety $X$, we have $\pi_{1}(X)\simeq\pi_{1}(X\times\mathbb{P}^{1})$. The requirement of $\varrho$ being big effectively excludes the case of uniruled varieties $X\times\mathbb{P}^{1}$, which is apparently non-hyperbolic.
In the next three sections, we will provide some applications of Theorem˜3.7.
4 On Campana’s abelianity conjecture
As our first application of Theorem˜3.7, in this section we prove Campana’s abelianity conjecture in the context of representations in positive characteristic.
4.1 Special and $h$-special varieties : properties and conjectures
We first recall the definition of special varieties by Campana [Cam04, Cam11].
Definition 4.1 (Campana’s specialness).
Let $X$ be a quasi-projective normal variety.
  1. (i)
    $X$ is weakly special if for any finite étale cover $\widehat{X}\to X$ and any proper birational modification $\widehat{X}^{\prime}\to\widehat{X}$, there exists no dominant morphism $\widehat{X}^{\prime}\to Y$ with connected general fibers such that $Y$ is a positive-dimensional quasi-projective variety of log general type.
  2. (ii)
    $X$ is special if for any proper birational modification $X^{\prime}\to X$ there is no dominant morphism $X^{\prime}\to Y$ with connected general fibers over a positive-dimensional quasi-projective variety $Y$ such that the Campana orbifold base (or simply orbifold base) is of log general type.
  3. (iii)
    $X$ is Brody special if it contains a Zariski dense entire curve.
Campana defined $X$ to be $H$-special if $X$ has vanishing Kobayashi pseudo-distance. Motivated by [Cam11, 11.3 (5)], in [CDY22, Definition 1.11] we introduce the following definition.
Definition 4.2 ($h$-special).
Let $X$ be a smooth quasi-projective variety. We define the equivalence relation $x\sim y$ of two points $x,y\in X$ iff there exists a sequence of holomorphic maps $f_{1},\ldots,f_{l}:\mathbb{C}\to X$ such that letting $Z_{i}\subset X$ to be the Zariski closure of $f_{i}(\mathbb{C})$, we have
$x\in Z_{1},Z_{1}\cap Z_{2}\not=\emptyset,\ldots,Z_{l-1}\cap Z_{l}\not=\emptyset,y\in Z_{l}.$
We set $R=\{(x,y)\in X\times X;x\sim y\}$. We define $X$ to be hyperbolically special ($h$-special for short) iff $R\subset X\times X$ is Zariski dense.
By definition, rationally connected projective varieties are $h$-special without referring to a theorem of Campana and Winkelmann [CW16], who proved that all rationally connected projective varieties contain Zariski dense entire curves. It also has the following properties.
Lemma 4.3 ([CDY22, Lemmas 10.2 & 10.3 & 10.4]).
  1. (a)
    If a smooth quasi-projective variety $X$ is Brody special, then it is $h$-special.
  2. (b)
    Let $X$ be an $h$-special quasi-projective variety. Let $S$ be a quasi-projective variety and let $p:X\to S$ be a dominant morphism. Then $S$ is $h$-special.
  3. (c)
    Let $X$ be an $h$-special smooth quasi-projective variety, and let $p:X^{\prime}\to X$ be a finite étale morphism or proper birational morphism from a quasi-projective variety $X^{\prime}$. Then $X^{\prime}$ is $h$-special. ∎
Proposition 4.4 ([CDY22, Proposition 11.14]).
If a quasi-projective smooth variety $X$ is special or $h$-special, the quasi-albanese map $a:X\to A$ of $X$ is $\pi_{1}$-exact, i.e., we have the following exact sequence :
$\pi_{1}(F)\to\pi_{1}(X)\to\pi_{1}(A)\to 1,$
where $F$ is a general fiber of $a$. ∎
In [Cam04, Cam11], Campana proposed the following tantalizing abelianity conjecture.
Conjecture 4.5 (Campana).
A special smooth projective variety has virtually abelian fundamental group.
In [CDY22] we observed that Conjecture˜4.5 fails for quasi-projective varieties. As illustrated in [CDY22, Example 11.26], we constructed a smooth quasi-projective variety such that it is both special and Brody special, yet it has nilpotent fundamental group that is not virtually abelian. Consequently, within the quasi-projective context, we revised Conjecture˜4.5 as follows.
Conjecture 4.6 ([CDY22, Conjecture 1.14]).
A special or $h$-special smooth quasi-projective variety has virtually nilpotent fundamental group.
In [CDY22], we confirm Conjecture˜4.6 for quasi-projective varieties with linear fundamental groups in characteristic zero.
Theorem 4.7 ([CDY22, Theorem E]).
Let $X$ be a special or $h$-special smooth quasiprojective variety. Let $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(\mathbb{C})$ be a linear representation. Then $\varrho(\pi_{1}(X))$ is virtually nilpotent. ∎
By [CDY22, Example 11.26], Theorem˜4.7 is shown to be sharp. In the context of representations in positive characteristic, we can obtain a stronger result as motivated by the following lemma.
Lemma 4.8.
Let $\Gamma\subset{\rm GL}_{N}(K)$ be a finitely generated subgroup where $K$ is an algebraically closed field of positive characteristic. If $\Gamma$ is virtually nilpotent, then it is virtually abelian.
Démonstration.
After replacing $\Gamma$ by a suitable finite index subgroup, we can assume that the Zariski closure $G$ of $\Gamma$ in ${\rm GL}_{N}$ is connected and nilpotent. Hence we have $\mathcal{D}G\subset R_{u}(G),$ where $R_{u}(G)$ is the unipotent radical of $G$ and $\mathcal{D}G$ is the the derived group of $G$. Consequently, We have
$\Gamma,\Gamma]\subset[G(K),G(K)]\subset R_{u}(G)(K).$
Note that $R_{u}(G)$ is unipotent. It is thus a successive extension of $\mathbb{G}_{a,K}$.
Since $\Gamma$ is finitely generated and nilpotent, it follows that $\Gamma,\Gamma]$ is also finitely generated by [ST00, Corollary 5.45]. From this fact we conclude that $\Gamma,\Gamma]$ is thus a successive extension of finitely generated subgroups of $\mathbb{G}_{a,K}$. Since finitely generated $p$-groups are finite, it implies that $\Gamma,\Gamma]$ is a finite group. Since $\Gamma$ is residually finite by Malcev’s theorem, it has a finite index normal subgroup $\Gamma_{1}$ such that $\Gamma_{1}\cap[\Gamma,\Gamma]=\{e\}$. Consequently, $\Gamma_{1},\Gamma_{1}]\subset[\Gamma,\Gamma]\cap\Gamma_{1}=\{e\}$. Hence $\Gamma_{1}$ is abelian. We conclude that $\Gamma$ is virtually abelian. ∎
4.2 A factorization result (I)
In this subsection we prove a factorization result.
Proposition 4.9.
Let $f:X\to Y$ be a dominant morphism from a smooth quasi-projective variety $X$ to a normal quasi-projective variety $Y$ such that general fibers of $f$ are connected. Let $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ where $K$ is any field. Assume that for a general smooth fiber $F$ of $f$, $\varrho({\rm Im}[\pi_{1}(F)\to\pi_{1}(X)])$ is finite. Then there exists
  • a generically finite proper surjective morphism $\mu:X_{1}\to X$ obtained by the composition of birational modifications and finite étale Galois covers ;
  • a dominant morphism $f_{1}:X_{1}\to Y_{1}$ onto a smooth quasi-projective variety $Y_{1}$ with connected general fibers ;
  • a generically finite dominant morphism $\nu:Y_{1}\to Y$ ;
  • a representation $\tau:\pi_{1}(Y_{1})\to{\rm GL}_{N}(K)$,
such that $f_{1}^{*}\tau=\mu^{*}\varrho$ and we have the following commutative diagram
[…]
Démonstration.
Step 1 : we can assume that $\varrho({\rm Im}[\pi_{1}(F)\to\pi_{1}(X)])$ is trivial. Let $Y^{\circ}$ be the Zariski open subset of $Y$ such that it is smooth and $f$ is smooth over $Y^{\circ}$. Denote by $X^{\circ}:=f^{-1}(Y^{\circ})$. We take a fiber $F:=f(y)$ with $y\in Y^{\circ}$, which is smooth and connected. Since $\Gamma:=\varrho(\pi_{1}(X))$ is residually finite by Malcev’s theorem, we can find a finite index normal subgroup $N\triangleleft\Gamma$ such that
$N\cap\varrho({\rm Im}[\pi_{1}(F)\to\pi_{1}(X)])=\{e\}.$
Let $\mu:X_{1}\to X$ be a finite étale cover such that
$\mu^{*}\varrho(\pi_{1}(X_{1}))=N.$
Let $X_{1}\stackrel{{\scriptstyle f_{1}}}{{\to}}Y_{1}\stackrel{{\scriptstyle\nu}}{{\to}}Y$ be the quasi-Stein factorization of $f\circ\mu$.
[…]
Then $f_{1}$ is smooth over $Y_{1}^{\circ}:=\nu^{-1}(Y^{\circ})$. Take any point $y_{1}\in\nu^{-1}(y)$. Then the fiber $F_{1}:=f^{-1}(y_{1})$ is smooth and
$\mu^{*}\varrho({\rm Im}[\pi_{1}(F_{1})\to\pi_{1}(X_{1})])\subset N\cap\varrho({\rm Im}[\pi_{1}(F)\to\pi_{1}(X)])=\{e\}$
as $\mu(F_{1})\subset F$.
In the following, to lighten the notation we will replace $X$, $Y$, $f$ and $\varrho$ by $X_{1}$, $Y_{1}$, $f_{1}$ and $\mu^{*}\varrho$ respectively, and assume that $\varrho({\rm Im}[\pi_{1}(F)\to\pi_{1}(X)])$ is trivial for a general smooth fiber $F$ of $f$.
Step 2. Compactifications and first reduction step. Let $\overline{X^{\prime}}\supset X$ and $\overline{Y}\supset Y$ be smooth projective compactifications of $X$ and $Y$, respectively, such that $f$ extends to a morphism $\bar{f}^{\prime}:\overline{X^{\prime}}\to\overline{Y}$. By performing a suitable blow-up of $\overline{X^{\prime}}$, we may assume that $\overline{X^{\prime}}$ is smooth. Since $f$ is dominant, it follows that $\bar{f}^{\prime}$ is surjective. Set $\overline{X}:=(\bar{f}^{\prime})^{-1}(Y)$. Then $\bar{f}:=\bar{f}^{\prime}|_{\overline{X}}$ is a surjective proper morphism whose general fibers are connected. As $Y$ is normal, it follows from Zariski’s Main Theorem that all fibers of $\bar{f}$ are connected.
Claim 4.10.
We may assume that $\bar{f}:\overline{X}\to Y$ is equidimensional.
Indeed, by the Hironaka-Gruson-Raynaud flattening theorem, there exists a birational proper morphism $Y_{1}\to Y$ from a smooth quasi-projective variety $Y_{1}$, such that for the irreducible component $T$ of $\overline{X}\times_{Y}Y_{1}$ dominating $Y_{1}$, the induced morphism $f_{T}:=T\to Y_{1}$ is surjective, proper, and flat.
[…]
Let $\nu:\overline{X}_{1}\to T$ denote the normalization map. Since $f_{T}$ is flat, the induced morphism $f_{1}:\overline{X}_{1}\to Y_{1}$ has equidimensional fibers. Let $\mu:\overline{X}_{1}\to\overline{X}$ be the induced proper birational morphism, and set $X_{1}:=\mu^{-1}(X)$.
Then one has a diagram
[…]
where the horizontal maps are proper birational, and the two spaces on the left satisfy the hypotheses of the proposition if we take the representation induced on $\pi_{1}(X_{1})$. Clearly, it suffices to show the result where $X$ (resp. $Y$) is replaced by $X_{1}$ (resp. $Y_{1}$). In the following, we may also replace $\overline{X}$ (resp. $Y$) by $\overline{X}_{1}$ and $Y$ (resp. $Y_{1}$).
Step 3. Induced representation on an open subset of $Y$. Consider a Zariski open set $Y^{\circ}\subset Y$ such that $X^{\circ}:=f^{-1}(Y^{\circ})$ is a topologically locally trivial fibration over $Y^{\circ}$. Recall that in Step 2, we arranged that $\YY\backslash Y^{\circ}$ is a simple normal crossing divisor. Then we have a short exact sequence
$\pi_{1}(F)\to\pi_{1}(X^{\circ})\to\pi_{1}(Y^{\circ})\to 1$
where $F$ is a general fiber of $f$ over $Y^{\circ}$. By Step 1, $\varrho({\rm Im}[\pi_{1}(F)\to\pi_{1}(X)])$ is trivial. Hence we can pass to the quotient, which yields a representation $\tau_{1}:\pi_{1}(Y^{\circ})\to{\rm GL}_{N}(K)$ so that $\varrho|_{\pi_{1}(X^{\circ})}=f^{\ast}\tau_{1}$.
Step 4. Reducing $Y$, we may assume that all divisorial components of $Y-Y^{\circ}$ intersect $f(X)$. Denote by $E$ the sum of prime divisors of $Y$ contained in the complement $\YY\backslash Y^{\circ}$. We decompose $E=E_{1}+E_{2}$ so that $E_{1}$ is the sum of prime divisors of $E$ that do not intersect $f(X)$. We replace $Y$ by $Y\backslash E_{1}$. Then for any prime divisor $P$ contained in $\YY\backslash Y^{\circ}$, $f^{-1}(P)\cap X\neq\varnothing$.
Step 5. Second use of Malcev’s theorem. Recall that $\YY\backslash Y^{\circ}$ is a simple normal crossing divisor $D:=\sum_{i\in I}D_{i}$ and $f^{-1}(D_{i})\neq\varnothing$ for each $i\in I$. Since $\bar{f}:\overline{X}\to Y$ is equidimensional, then for any prime component $P$ of $f^{-1}(D_{i})$, the morphism $f|_{P}:P\to D_{i}$ is dominant. Also, since $X$ is normal, $X$ is smooth at the general points of $P$.
This allows to find a smooth point $x$ in $P$ (resp. $y\in D_{i}\backslash\bigcup_{j\in I;j\neq i}D_{j}$) with local coordinates $(z_{1},\dotsc z_{m})$ (resp. $(w_{1},\dotsc,w_{n})$) around $x$ (resp. $y$), such that around $x$ (resp. around $y$) we have $P=(z_{1}=0)$ (resp. $D=(w_{1}=0)$) and $f^{\ast}(w_{1})=z_{1}^{k}$ for some $k\geq 1$. Hence the small meridian loop $\gamma$ around the general point of $P$ is mapped to $\eta_{i}^{k}$ where $\eta_{i}$ is the small meridian loop around $D_{i}$. On the other hand, since $\gamma$ is trivial in $\pi_{1}(X)$, it follows that
$0=\varrho(\gamma)=\tau_{1}(\eta_{i}^{k}).$ (4.1)
Hence $\tau_{1}(\eta_{i})$ is a torsion element.
[…]
$D\cap U=(\cup_{i=1}^{k}D_{j_{i}})\cap U$
with $D_{j_{i}}=(z_{i}=0)$ for each $i$. Set $U^{*}:=U-D$ and we fix a point $y_{J}\in U^{*}$. The fundamental group $\pi_{1}(U^{*},y_{J})\simeq\mathbb{Z}^{k}$ is abelian. Let $e_{i}$ be a clockwise loop with basepoint $y_{J}$ around the origin in the $i$-th factor of $(\mathbb{D}^{*})^{k}\times\mathbb{D}^{n-k}$. Pick a smooth path $h:[0,1]\to Y^{\circ}$ connecting $y_{0}$ with $y_{J}$, and denote by $\gamma_{i}\in\pi_{1}(Y^{\circ},y_{0})$ the homotopy class of the loop $h^{-1}\cdot e_{i}\cdot h$. Set $T_{i}:=\tau_{1}(\gamma_{i})$. Clearly, $T_{1},\ldots,T_{k}$ commute pairwise. By (4.1), we know that each $T_{i}$ is an torsion element. Therefore,
$G_{J,i}:=\left\{T_{1}^{\ell_{1}}\cdots T_{k}^{\ell_{k}}\mid\ell_{1},\ldots,\ell_{k}\in\mathbb{Z}\right\}$ (4.2)
is a finite abelian subgroup of $\tau_{1}(\pi_{1}(Y^{\circ},y_{0}))$. Note that $G_{J,i}$ depends on the choice of $h$, and the above coordinate system $(U;z_{1},\ldots,z_{k})$, but any two such groups are conjugate.
By Malcev’s theorem, $\tau_{1}(\pi_{1}(Y^{\circ},y_{0}))$ is a finitely generated, residually finite group. Hence there exists a finite-index normal subgroup $\Gamma$ of $\tau_{1}(\pi_{1}(Y^{\circ},y_{0}))=\varrho(\pi_{1}(X,x_{0}))$ such that
$\Gamma\cap G_{J,i}=\{e\}$ (4.3)
for each $J\in\mathcal{S}$ and $i\in\{1,\ldots,k_{J}\}$. Let $\mu:X_{1}\to X$ be the finite Galois étale cover such that $\mu^{*}\varrho(\pi_{1}(X_{1}))=\Gamma$.
Step 6. Completion of the proof. Let ${X}_{1}\stackrel{{\scriptstyle{f}_{1}}}{{\to}}Y_{1}\stackrel{{\scriptstyle\nu}}{{\to}}Y$ be the quasi-Stein factorization of ${f}\circ\mu$. Then $f_{1}$ is smooth over $Y_{1}^{\circ}:=\nu^{-1}(Y^{\circ})$. We take blow-ups $\mu_{2}:X_{2}\to X_{1}$ and $\nu_{2}:Y_{2}\to Y_{1}$ such that
  • both $X_{2}$ and $Y_{2}$ are smooth and $f_{1}$ induces a morphism $f_{2}:X_{2}\to Y_{2}$.
  • The map $Y_{2}^{\circ}:=\nu^{-1}_{2}(Y_{1}^{\circ})\to Y_{1}^{\circ}$ is an isomorphism.
[…]
  • Set $X_{2}^{\circ}:=f_{2}^{-1}(Y_{2}^{\circ})$. Then $f_{2}^{\circ}:=f_{2}|_{X_{2}^{\circ}}$ is a topologically locally trivial fibration over $Y_{2}^{\circ}$.
Note that for the representation $(\nu_{2}\circ\nu)^{*}\tau_{1}:\pi_{1}(Y_{2}^{\circ})\to{\rm GL}_{N}(K)$, its image
$(\nu\circ\nu_{2})^{*}\tau_{1}(\pi_{1}(Y_{2}^{\circ}))=(\mu\circ\mu_{2})^{*}\varrho(\pi_{1}(X_{2}))\subset\Gamma.$
Let $a:\mathbb{D}\to Y_{2}$ be any holomorphic disk such that $(\nu\circ\nu_{2}\circ a)^{-1}(D)=\{0\}$. It follows that $\nu\circ\nu_{2}\circ a(0)\in D_{J,i}$ for some $J\in\mathcal{S}$ and $i\in\{1,\ldots,k_{J}\}$. Then $(\nu\circ a)^{*}\tau_{1}(\pi_{1}(\mathbb{D}^{*}))$ is contained in some conjugacy class of $G_{J,i}$. Since $\Gamma\cap G_{J,i}=\{e\}$ by (4.3) and given that $\Gamma$ is a normal subgroup of $\nu^{*}\tau_{1}(\pi_{1}(Y_{1}^{\circ}))$, it implies that $(\nu\circ a)^{*}\tau_{1}(\pi_{1}(\mathbb{D}^{*}))=\{e\}$. Consequently, the pullback $(\nu_{2}\circ\nu)^{*}\tau_{1}$ extends to a representation $\tau:\pi_{1}(Y_{2})\to{\rm GL}_{N}(K)$. It follows that
$f_{2}^{*}\tau=(\mu\circ\mu_{2})^{*}\varrho.$
We now change the notation by renaming $X_{2}$ as $X_{1}$, $Y_{2}$ as $Y_{1}$, $f_{2}$ as $f_{1}$, $\mu\circ\mu_{2}$ as $\mu$, and $\nu\circ\nu_{2}$ as $\nu$. This completes the proof of the proposition. ∎
Remark 4.11.
Proposition˜4.9 is proven in [CDY22, Proposition 2.5] in cases where ${\rm char}\,K=0$ and the proof utilizes Selberg’s lemma : a finitely generated linear group in characteristic zero is virtually torsion free.
4.3 On the Abelianity conjecture
Theorem 4.12.
Let $X$ be a smooth quasi-projective variety, and let $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be any representation with $K$ any field of positive characteristic. If $X$ is special or $h$-special, then $\varrho(\pi_{1}(X))$ is virtually abelian.
Démonstration.
Step 1. We prove that $\varrho(\pi_{1}(X))$ is solvable. We may assume that $K$ is algebraically closed. Let $G$ be the Zariski closure of $\varrho(\pi_{1}(X))$. Note that any finite étale cover of a special (resp. $h$-special) variety is still special (resp. $h$-special). After replacing $X$ by a finite étale cover, we may assume that $G$ is connected. Let $R(G)$ be the radical of $G$. Let $H:=G/R(G)$, which is semisimple. If $\dim H>0$, then $\varrho$ induces a Zariski dense representation $\sigma:\pi_{1}(X)\to H(K)$. Let ${\rm sh}_{\sigma}:X\to{\rm Sh}_{\sigma}(X)$ be the Shafarevich morphism of $\sigma$. By the property of the Shafarevich morphism in Theorem˜2.9, a general fiber $F$ of ${\rm sh}_{\sigma}$ is connected and $\sigma({\rm Im}[\pi_{1}(F)\to\pi_{1}(X)])$ is finite. We apply Proposition˜4.9 to conclude that there exist
  1. (i)
    a generically finite proper surjective morphism $\mu:X_{1}\to X$ from a smooth quasi-projective variety obtained by the composition of birational modifications and finite étale Galois covers ;
  2. (ii)
    a generically finite dominant morphism $\nu:Y_{1}\to{\rm Sh}_{\sigma}(X)$ ;
  3. (iii)
    a dominant morphism $f_{1}:X_{1}\to Y_{1}$ with $Y_{1}$ a smooth quasi-projective variety with connected general fibers ;
  4. (iv)
    a representation $\tau:\pi_{1}(Y_{1})\to H(K)$
such that we have following commutative diagram
[…]
and $\mu^{*}\sigma=f_{1}^{*}\tau$. We can show that $\tau$ is a big representation. Thanks to Theorem˜3.7, $Y_{1}$ is of log general type and pseudo Picard hyperbolic. This leads to a contradiction since $X$ is special (thus weakly special by [Cam11]) or $h$-special. Hence $G=R(G)$.
Step 2. We prove that $\varrho(\pi_{1}(X))$ is virtually abelian. Note that any finite étale cover of a special (resp. $h$-special) variety is still special (resp. $h$-special) by [Cam04] and Lemma˜4.3. Replacing $X$ by a finite étale cover, we may assume that $\pi_{1}(X)^{ab}\to\pi_{1}(A)$ is an isomorphism, where $\pi_{1}(X)^{ab}:=\pi_{1}(X)/[\pi_{1}(X),\pi_{1}(X)]$ and $A$ is the quasi-albanese variety. Since $X$ is special or $h$-special, by Proposition˜4.4, the quasi-albanese map $a:X\to A$ of $X$ is $\pi_{1}$-exact, i.e., we have the following exact sequence :
$\pi_{1}(F)\to\pi_{1}(X)\to\pi_{1}(A)\to 1,$
where $F$ is a general fiber of $a$. Hence $\pi_{1}(X),\pi_{1}(X)]$ is the image of $\pi_{1}(F)\to\pi_{1}(X)$, which is thus finitely generated. It implies that $\varrho(\pi_{1}(X)),\varrho(\pi_{1}(X))]=\varrho([\pi_{1}(X),\pi_{1}(X)])$ is also finitely generated. By Step 1, $G$ is solvable. Hence we have $\mathcal{D}G\subset R_{u}(G),$ where $R_{u}(G)$ is the unipotent radical of $G$ and $\mathcal{D}G$ is the the derived group of $G$. Consequently, we have
$\varrho(\pi_{1}(X)),\varrho(\pi_{1}(X))]\subset[G(K),G(K)]\subset R_{u}(G)(K).$
Note that every subgroup of finite index in $\pi_{1}(X),\pi_{1}(X)]$ is also finitely generated (cf. [ST00, Proposition 4.17]. By the same arguments in Lemma˜4.8, we conclude that $\varrho(\pi_{1}(X)),\varrho(\pi_{1}(X))]$ is finite. Hence $\varrho(\pi_{1}(X))$ is virtually abelian.
4.4 A characterization of semiabelian varieties.
Theorem˜4.12 allows us to give a characterization of semiabelian varieties.
Proposition 4.13.
Let $Y$ be a smooth quasi-projective variety, and let $\varrho:\pi_{1}(Y)\to{\rm GL}_{N}(K)$ be a big representation where $K$ is a field of positive characteristic.
  1. (i)
    If $Y$ is special or $h$-special, then there exists a finite étale cover $X$ of $Y$, such that its quasi-Albanese map $\alpha:X\to A$ is birational and $\alpha_{*}:\pi_{1}(X)\to\pi_{1}(A)$ is an isomorphism.
  2. (ii)
[…]
The proof closely follows that of [CDY22, Proposition 12.7], and we present it here for the sake of completeness.
Démonstration.
Proof of (i). By Theorem˜4.12, there is a finite étale cover $X$ of $Y$ such that $G:=\varrho(\pi_{1}(X))$ is abelian and torsion free. It follows that $\varrho|_{\pi_{1}(X)}:\pi_{1}(X)\to{\rm GL}_{N}(K)$ factors through $H_{1}(X,\mathbb{Z})/{\rm torsion}$. By [CDY22, Lemma 11.5], $\alpha$ is dominant with connected general fibers. Since $\alpha_{*}:H_{1}(X,\mathbb{Z})/{\rm torsion}\to H_{1}(A,\mathbb{Z})$ is isomorphic, $\varrho$ further factors through $H_{1}(A,\mathbb{Z})$.
[…]
From the above diagram for every fiber $F$ of $\alpha$, $\varrho({\rm Im}[\pi_{1}(F)\to\pi_{1}(X)])$ is trivial. Since $\varrho$ is big, the general fiber of $\alpha$ is thus a point. Hence $\alpha$ is birational. Since $\alpha:X\to A$ is $\pi_{1}$-exact by Proposition˜4.4, it follows that $\alpha_{*}:\pi_{1}(X)\to\pi_{1}(A)$ is an isomorphism.
Proof of (ii). If $\bar{\kappa}(Y)=0$, then $Y$ is special by [Cam11, Corollary 5.6] and for any finite étale cover $X$ of $Y$, we have $\bar{\kappa}(X)=0$ by [CDY22, Lemma 6.7]. By ˜4.13.(i), there is a finite étale cover $X$ of $Y$ such that its quasi-Albanese map $\alpha:X\to A$ is birational. We apply [CDY22, Lemma 3.2] to conclude that $\alpha$ is proper in codimension one. ∎
5 A structure theorem : on a conjecture by Kollár
In [Kol95, Conjecture 4.18], Kollár raised the following conjecture on the structure of varieties with big fundamental group.
Conjecture 5.1.
Let $X$ be a smooth projective variety with big fundamental group such that $0<\kappa(X)<\dim X$. Then $X$ has a finite étale cover $p:X^{\prime}\to X$ such that $X^{\prime}$ is birational to a smooth family of abelian varieties over a projective variety of general type $Z$ which has big fundamental group.
In this section we address Conjecture˜5.1. Our theorem is the following :
Theorem 5.2.
Let $X$ be a quasi-projective normal variety and let $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be a big representation where $K$ is a field of positive characteristic. Then
  1. (i)
    the logarithmic Kodaira dimension satsifies $\bar{\kappa}(X)\geq 0$.
  2. (ii)
    There is a proper Zariski closed subset $\Xi$ of $X$ such that each non-constant algebraic morphism $\mathbb{A}^{1}\to X$ has image in $\Xi$.
  3. (iii)
    If $0<\bar{\kappa}(X)<\dim X$, after replacing $X$ by a finite étale cover and a birational modification, there are a semi-abelian variety $A$, a quasi-projective manifold $V$ and a birational morphism $a:X\to V$ such that the following commutative diagram holds :
[…]
    where $j$ is the logarithmic Iitaka fibration of $X$ and $h:V\to J(X)$ is a locally trivial fibration with fibers isomorphic to $A$. Moreover, for a general fiber $F$ of $j$, $a|_{F}:F\to A$ is proper in codimension one.
In our previous work [CDY22, Theorem 12.5], we established Theorem˜5.2 for the cases where ${\rm char}\,K=0$ and $\varrho$ is big and reductive. The proof of Theorem˜5.2 closely follows that of [CDY22, Theorem 12.5]. Note that ˜4.13.(ii) corresponds to the case $\bar{\kappa}(X)=0$ of ˜5.2.(iii).
Démonstration.
We may assume that $K$ is algebraically closed. To prove the theorem we are free to replace $X$ by a birational modification and by a finite étale cover since the logarithmic Kodaira dimension will remain unchanged. We replace $\varrho$ by its semisimplification, which is still big by Lemma˜2.1. Hence we might assume that $\varrho$ is big and semisimple. Consequently, after replacing $X$ by a finite étale cover, the Zariski closure $G$ of $\varrho$ is reductive and connected. Let $\mathcal{D}G$ be the derived group of $G$, which is semisimple. Let $Z\subset G$ be the maximal central torus of $G$. Then $T:=G/\mathcal{D}G$ is a torus and $S:=G/Z$ is semisimple. The natural morphism $G\to S\times T$ is a central isogeny (see [Con14, Corollary 5.3.3]). The induced representation $\varrho^{\prime}:\pi_{1}(X)\to S(K)\times T(K)$ by $\varrho$ is also big. Consider the representation $\sigma:\pi_{1}(X)\to S(K)$, obtained by composing $\varrho$ with the morphism $G\to S$. Then $\sigma(\pi_{1}(X))$ is Zariski dense. Let ${\rm sh}_{\sigma}:X\to{\rm Sh}_{\sigma}(X)$ be the Shafarevich morphism of $\sigma$.
Like Step 1 of the proof of Theorem˜4.12, we apply Proposition˜4.9 to conclude that there exist
  1. (i)
    a generically finite proper surjective morphism $\mu:X_{1}\to X$ from a smooth quasi-projective variety obtained by the composition of birational modifications and finite étale Galois covers ;
  2. (ii)
    a generically finite dominant morphism $\nu:Y_{1}\to{\rm Sh}_{\sigma}(X)$ ;
  3. (iii)
    a dominant morphism $f_{1}:X_{1}\to Y_{1}$ with $Y_{1}$ a smooth quasi-projective variety with connected general fibers ;
  4. (iv)
    a big representation $\tau:\pi_{1}(Y_{1})\to S(K)$ ;
such that we have the following commutative diagram
[…]
and $\mu^{*}\sigma=f_{1}^{*}\tau$. It is straightforward to show that $\tau$ is a big representation. Thanks to Theorem˜3.7, the special loci $\mathrm{Sp}_{\mathrm{alg}}(Y_{1})$ and $\mathrm{Sp}_{\mathrm{p}}(Y_{1})$ are both proper Zariski closed subset of $Y_{1}$. In particular, $Y_{1}$ is of log general type. Note that $Y_{1}$ can be a point.
Consider the morphism
$g:X_{1}$ $\to A\times Y_{1}$
$x$ $\mapsto(\alpha(x),f_{1}(x)).$
where $\alpha:X_{1}\to A$ is the quasi Albanese map of $X_{1}$.
Claim 5.3.
We have $\dim X_{1}=\dim g(X_{1})$.
Démonstration.
For a general smooth fiber $F$ of $f_{1}$, $\mu^{*}\sigma({\rm Im}[\pi_{1}(F)\to\pi_{1}(X)])$ is trivial. Since $\mu^{*}\varrho^{\prime}:\pi_{1}(X_{1})\to S(K)\times T(K)$ is big, by the construction of $\sigma$, we conclude that the representation $\eta:\pi_{1}(F)\to T(K)$ obtained by
$\pi_{1}(F)\to\pi_{1}(X_{1})\stackrel{{\scriptstyle\mu^{*}\varrho^{\prime}}}{{\to}}S(K)\times T(K)\to T(K)$
is big. Since $T(K)$ is commutative, similar to (4.4), $\eta$ factors through $\pi_{1}(F)\to\pi_{1}(A)\to T(K)$. This implies that $\dim F=\dim\alpha(F)$. Hence $\dim X_{1}=\dim g(X_{1})$. ∎
Let us prove ˜5.2.(i). Thanks to Claim˜5.3, for a general smooth fiber $F$ of $f_{1}$, we have $\dim F=\dim\alpha(F)$. Hence $\bar{\kappa}(F)\geq 0$. Since $Y_{1}$ is of log general type, by the subadditivity of the logarithmic Kodaira dimension proven in [Fuj17, Theorem 1.9], we obtain
$\bar{\kappa}(X_{1})\geq\bar{\kappa}(Y_{1})+\bar{\kappa}(F)\geq\bar{\kappa}(Y_{1})=\dim Y_{1}\geq 0.$
Hence $\bar{\kappa}(X)=\bar{\kappa}(X_{1})\geq 0$. The first claim is proved.
Let us prove ˜5.2.(ii). By Lemma˜3.4, the special subset ${\rm Sp}(\varrho)$ defined in Definition˜0.1 is a proper closed subset of $X$. Let $\gamma:\mathbb{A}^{1}\to X$ be a non-constant algebraic morphism. Then $\gamma^{*}\varrho(\pi_{1}(\mathbb{A}^{1}))=\{1\}$. By the definition of ${\rm Sp}(\varrho)$, we have $\gamma(\mathbb{A}^{1})\subset{\rm Sp}(\varrho)$.
Finally ˜5.2.(iii) follows from [CDY22, Theorem 12.1]. ∎
6 On the Shafarevich conjecture and a conjecture by Claudon-Höring-Kollár
In this section, we will work on projective varieties.
6.1 Algebraic varieties with compactifiable universal cover
In the paper [CHK13, CH13], Claudon, Höring and Kollár proposed the following intriguing conjecture :
Conjecture 6.1.
Let $X$ be a complex projective manifold with infinite fundamental group $\pi_{1}(X)$. Suppose that the universal cover $\widetilde{X}$ is quasi-projective. Then after replacing $X$ by a finite étale cover, there exists a locally trivial fibration $X\rightarrow A$ with simply connected fiber $F$ onto a complex torus $A$. In particular we have $\widetilde{X}\simeq F\times\mathbb{C}^{\operatorname{dim}A}$.
It’s worth noting that assuming abundance conjecture, Claudon, Höring and Kollár proved this conjecture in [CHK13], thereby providing an unconditional proof for Conjecture˜6.1 in cases where $\dim X\leqslant 3$. Additionally, Claudon and Höring, in [CH13], proved Conjecture˜6.1 when $\pi_{1}(X)$ is virtually abelian, a result essential for the proof of Corollary˜6.2.
As an application of Theorem˜4.12, we establish a linear version of Conjecture˜6.1 without relying on the abundance conjecture.
Corollary 6.2.
Let $X$ be a smooth projective variety with an infinite fundamental group $\pi_{1}(X)$, such that its universal covering $\widetilde{X}$ is a Zariski open subset of some compact Kähler manifold $\overline{X}$. If there exists a faithful representation $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K),$ where $K$ is any field, then the Albanese map of $X$ is (up to finite étale cover) locally isotrivial with simply connected fiber $F$. In particular we have $\widetilde{X}\simeq F\times\mathbb{C}^{q(X)}$ with $q(X)$ the irregularity of $X$.
Démonstration.
The proof is suggested by a referee of our paper. Consider the core map
$c_{X}:X\dashrightarrow Y:=C(X)$
defined by Campana, such that the orbifold base of $c_{X}$ is of orbifold general type (see [Cam04]). By the orbifold version of the Kobayashi-Ochiai theorem [Cam04, Theorem 8.2], the composed meromorphic map
$\widetilde{X}\xrightarrow{\pi}X\overset{c_{X}}{\dashrightarrow}Y$
extends to a meromorphic map from a Kähler compactification $\overline{\widetilde{X}}$ of $\widetilde{X}$. This implies that the general fiber $F$ of $c_{X}$ satisfies that $\pi^{-1}(F)$ has only finitely many connected components. In particular, the induced homomorphism $\pi_{1}(F)\rightarrow\pi_{1}(X)$ has image of finite index.
According to [Cam04], $F$ is a special manifold. Since we assume the existence of a faithful representation $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$, it follows from Theorem˜F when $\mathrm{char}\,K>0$ and [Cam04, Theorem 7.8] when $\mathrm{char}\,K=0$ that the image $\varrho([\pi_{1}(F)\to\pi_{1}(X)])$ is virtually abelian. As the image of $\pi_{1}(F)\to\pi_{1}(X)$ has finite index, $\pi_{1}(X)$ itself is virtually abelian. The conclusion then follows from [CH13, Theorem 1.5], completing the proof of Corollary˜6.2. ∎
Remark 6.3.
As mentioned above, the proof of Corollary˜6.2 was suggested by a referee, following the original ideas of [CHK13] and [CH10, p. 7]. Our original proof was significantly more involved and relied on the hyperbolicity arguments in Theorem˜3.7.
6.2 A factorization result (II)
In this subsection, we will prove that any linear representation factors through its Shafarevich morphism after passing to a finite étale cover. It will be used in the proof of Theorem˜6.23.
Lemma 6.4.
Let $X$ be a projective normal variety. Let $F\subset X$ be a connected Zariski closed subset. Then there exists a connected open neighbourhood $U\subset X$ of $F$ such that ${\rm Im}[\pi_{1}(F)\to\pi_{1}(X)]={\rm Im}[\pi_{1}(U)\to\pi_{1}(X)]$.
Remark 6.5.
A stronger result holds as a consequence of [Hof09, Theorem 4.5]. Namely there exists a connected open neighbourhood $U\subset X$ of $F$ such that the induced map $\pi_{1}(F)\to\pi_{1}(U)$ is an isomorphism. Here we give a direct proof of Lemma˜6.4 for the sake of convenience.
Proof of Lemma˜6.4.
The construction of $U$ is as follows. For each $t\in F$, we take a connected open neighbourhood $\Omega_{t}\subset X$ of $t$ such that
  1. (a)
    any loop in $\Omega_{t}$ is null homotopic in $X$,
  2. (b)
    $F\cap\Omega_{t}$ is connected.
We take a connected open neighbourhood $W_{t}$ of $t$ such that $\overline{W_{t}}\subset\Omega_{t}$. Since $X$ is projective, $F$ is compact. We may take $t_{1},\ldots,t_{l}\in F$ such that $F\subset W_{t_{1}}\cup\cdots\cup W_{t_{l}}$. For each $s\in F$, we take a connected open neighbourhood $U_{s}$ of $s$ such that for each $i=1,\ldots,l$, if $s\in W_{t_{i}}$, then $U_{s}\subset W_{t_{i}}$, if $s\in\overline{W_{t_{i}}}$ then $U_{s}\subset\Omega_{t_{i}}$, and if $s\not\in\overline{W_{t_{i}}}$, then $U_{s}\cap\overline{W_{t_{i}}}=\emptyset$. We set $U=\cup_{s\in F}U_{s}$. Note that $U$ is open and $F\subset U$. Since $F$ and $U_{s}$ are connected, $U$ is connected. We shall show that $U$ satisfies the property of our lemma.
Let $\pi:\widetilde{X}\to X$ be the universal covering space. Set $\Gamma=\pi_{1}(X)$. The action $\Gamma\curvearrowright\widetilde{X}$ is free. We fix a connected component $F^{\prime}\subset\widetilde{X}$ of $\pi^{-1}(F)$. Define $H\subset\Gamma$ by $\gamma\in H$ iff $\gamma F^{\prime}=F^{\prime}$. Note that $F\subset\widetilde{X}$ is an analytic set, hence locally path connected. Hence $F^{\prime}\subset\widetilde{X}$ is path connected analytic set. Thus $H={\rm Im}[\pi_{1}(F)\to\pi_{1}(X)]$.
For each $t\in F$, we take $t^{\prime}\in F^{\prime}$ such that $t=\pi(t^{\prime})$. Then $F^{\prime}\cap\pi^{-1}(t)=\{\gamma t^{\prime};\gamma\in H\}$. For each $i=1,\ldots,l$, we denote by $\Omega_{t_{i}^{\prime}}^{\prime}\subset\widetilde{X}$ the connected component of $\pi^{-1}(\Omega_{t_{i}})$ which contains $t_{i}^{\prime}$. Since any loop in $\Omega_{t_{i}}$ is null homotopic in $X$, the natural map $\Omega_{t_{i}^{\prime}}^{\prime}\to\Omega_{t_{i}}$ is isomorphic. Then $\pi^{-1}(\Omega_{t_{i}})=\sqcup_{\gamma\in\Gamma}\gamma\Omega_{t_{i}^{\prime}}^{\prime}$ is disjoint union. Since $F\cap\Omega_{t_{i}}$ is connected, $\Omega_{t_{i}^{\prime}}^{\prime}$ does not intersect with $\gamma F^{\prime}$ for all $\gamma\in\Gamma-H$. Thus
$F^{\prime}\cap(\cup_{\gamma\in\Gamma-H}\gamma\Omega_{t_{i}^{\prime}}^{\prime})=\emptyset.$ (6.1)
Let $W_{t_{i}^{\prime}}^{\prime}\subset\Omega_{t_{i}^{\prime}}^{\prime}$ be the inverse image of $W_{t_{i}}$ under $\Omega_{t_{i}^{\prime}}^{\prime}\to\Omega_{t_{i}}$. Then $\pi^{-1}(\overline{W_{t_{i}}})=\cup_{\gamma\in\Gamma}\gamma\overline{W_{t_{i}^{\prime}}^{\prime}}$, which is a closed subset of $\widetilde{X}$. Set $\mathcal{C}_{i}=\cup_{\gamma\in\Gamma-H}\gamma\overline{W_{t_{i}^{\prime}}^{\prime}}$. By $\mathcal{C}_{i}=\pi^{-1}(\overline{W_{t_{i}}})\cap(\cup_{\gamma\in H}\gamma\Omega_{t_{i}^{\prime}}^{\prime})^{c}$, we conclude that $\mathcal{C}_{i}$ is a closed set. By (6.1), we have $F^{\prime}\cap\mathcal{C}_{i}=\emptyset$. We set $\mathcal{C}=\mathcal{C}_{1}\cup\cdots\cup\mathcal{C}_{l}$. Then $\mathcal{C}\subset\widetilde{X}$ is a closed subset and $F^{\prime}\cap\mathcal{C}=\emptyset$.
Now for each $t\in F$, we denote by $U^{\prime}_{t^{\prime}}\subset\widetilde{X}$ the connected component of $\pi^{-1}(U_{t})$ which contains $t^{\prime}$. We claim that if $\tau\in H$, then $\tau U^{\prime}_{t^{\prime}}\cap\mathcal{C}=\emptyset$, and if $\tau\in\Gamma-H$, then $\tau U^{\prime}_{t^{\prime}}\subset\mathrm{Int}(\mathcal{C})$, where $\mathrm{Int}(\mathcal{C})$ is the interior of $\mathcal{C}$. To prove this, we first suppose $\tau\in H$. To show $\tau U^{\prime}_{t^{\prime}}\cap\mathcal{C}=\emptyset$, it is enough to show that $\tau U^{\prime}_{t^{\prime}}\cap\gamma\overline{W_{t^{\prime}_{i}}^{\prime}}=\emptyset$ for each $\gamma\in\Gamma-H$ and $i=1,\ldots,l$. Indeed, if $t\in\overline{W_{t_{i}}}$, then $U_{t}\subset\Omega_{t_{i}}$. By $\gamma\not=\tau$, we have $\gamma\Omega^{\prime}_{t^{\prime}_{i}}\cap\tau\Omega^{\prime}_{t^{\prime}_{i}}=\emptyset$. Hence $\tau U^{\prime}_{t^{\prime}}\cap\gamma\overline{W_{t^{\prime}_{i}}^{\prime}}=\emptyset$. If $t\not\in\overline{W_{t_{i}}}$, then $U_{t}\cap\overline{W_{t_{i}}}=\emptyset$. Thus $\tau U^{\prime}_{t^{\prime}}\cap\gamma\overline{W_{t^{\prime}_{i}}^{\prime}}=\emptyset$. We have proved that $\tau U^{\prime}_{t^{\prime}}\cap\mathcal{C}=\emptyset$, if $\tau\in H$. Next suppose $\tau\in\Gamma-H$. We may take $i$ such that $t\in W_{t_{i}}$. Then by $U_{t}\subset W_{t_{i}}$, we have $\tau U^{\prime}_{t^{\prime}}\subset\tau W^{\prime}_{t^{\prime}_{i}}\subset\mathrm{Int}(\mathcal{C})$.
Now we set $U^{\prime}=\cup_{t\in F}\cup_{\gamma\in H}\gamma U^{\prime}_{t^{\prime}}$. Then we have $F^{\prime}\subset U^{\prime}$. Since $F^{\prime}$ and $\gamma U^{\prime}_{t^{\prime}}$ are connected, $U^{\prime}$ is connected. Note that $U^{\prime}\cap\mathcal{C}=\emptyset$ and $\gamma U^{\prime}=U^{\prime}$ for all $\gamma\in H$. On the other hand, for all $\gamma\in\Gamma-H$, we have $\gamma U^{\prime}\subset\mathrm{Int}(\mathcal{C})$. By $\pi^{-1}(U)=\cup_{\gamma\in\Gamma}\gamma U^{\prime}$, we conclude that $U^{\prime}$ is a connected component of $\pi^{-1}(U)$ and $\varrho({\rm Im}[\pi_{1}(U)\to\pi_{1}(X)])=H$. The proof is completed. ∎
Lemma 6.6.
Let $f:X\to S$ be a proper surjective map between quasi-projective varieties $X$ and $S$ such that all fibers of $f$ are connected. Let $\varrho:\pi_{1}(X)\to\Gamma$ be a morphism to a group $\Gamma$. Assume that for every fiber $F$ of $f$, the image $\varrho({\rm Im}[\pi_{1}(F)\to\pi_{1}(X)])$ is trivial. Then there exists a morphism $\tau:\pi_{1}(S)\to\Gamma$ such that $\tau\circ f_{*}=\varrho$.
Démonstration.
Let $p:X^{\prime}\to X$ be a Galois covering corresponding to $\mathrm{ker}(\varrho)\subset\pi_{1}(X)$. Replacing $\Gamma$ by $\varrho(\pi_{1}(X))$, we may assume that $\varrho$ is surjective. Then $\mathrm{Gal}(X^{\prime}/X)=\Gamma$. The action $\Gamma\curvearrowright X^{\prime}$ is properly discontinuous. Let $g:X^{\prime}\to S$ be the composite of $p:X^{\prime}\to X$ and $f:X\to S$. We define an equivalence relation $\sim$ on $X^{\prime}$ by $x\sim y$ iff $g(x)=g(y)$ and $y$ is contained in the connected component of $g^{-1}(g(x))$ which contains $x$. Let $\varphi:X^{\prime}\to S^{\prime}$ be the quotient map endowed with the quotient topology on $S^{\prime}$. Then we have the induced map $q:S^{\prime}\to S$, which is continuous. Each $\gamma\in\Gamma$ preserves the equivalence relation $\sim$ on $X^{\prime}$, namely if $x\sim y$, then $\gamma x\sim\gamma y$. Hence the action $\Gamma\curvearrowright X^{\prime}$ descends to $\Gamma\curvearrowright S^{\prime}$ as a continuous action.
We shall prove that $q:S^{\prime}\to S$ is a topological Galois covering with $\mathrm{Gal}(S^{\prime}/S)=\Gamma$. We take $s\in S$. We take a connected open neighborhood $U_{s}\subset X$ of $f^{-1}(s)$ as in Lemma˜6.4. Then
$\varrho({\rm Im}[\pi_{1}(U_{s})\to\pi_{1}(X)])=\{1\}.$ (6.2)
Since $f:X\to S$ is proper, we may take an open neighborhood $V_{s}\subset S$ of $s\in S$ such that $f^{-1}(V_{s})\subset U_{s}$. We replace $U_{s}$ by $f^{-1}(V_{s})$. Let $W\subset X^{\prime}$ be a connected component of $p^{-1}(U_{s})$. Then by (6.2), the induced map $p|_{W}:W\to U_{s}$ is a homeomorphism. We have a disjoint union
$p^{-1}(U_{s})=\bigcup_{\gamma\in\Gamma}\gamma W.$ (6.3)
Set $\Omega=\varphi(W)$. Then $\varphi^{-1}(\Omega)=W$, hence $\Omega\subset S^{\prime}$ is open. Since every fiber of $f:X\to S$ is connected, the induced map $q|_{\Omega}:\Omega\to V_{s}$ is injective. Since $f:X\to S$ is surjective, $q|_{\Omega}:\Omega\to V_{s}$ is surjective, hence bijective. Since $f:X\to S$ is proper, $(q|_{\Omega})^{-1}:V_{s}\to\Omega$ is continuous, hence $q|_{\Omega}:\Omega\to V_{s}$ is homeomorphic. By (6.3), we have a disjoint union $q^{-1}(V_{s})=\bigcup_{\gamma\in\Gamma}\gamma\Omega$. Hence $q:S^{\prime}\to S$ is a topological Galois covering with $\mathrm{Gal}(S^{\prime}/S)=\Gamma$. Thus we get $\tau:\pi_{1}(S)\to\Gamma$ such that $\tau\circ f_{*}=\varrho$. ∎
Lemma 6.7.
Let $f:X\to S$ be a proper surjective map between projective normal varieties $X$ and $S$ such that all fibers of $f$ are connected. Let $\varrho:\pi_{1}(X)\to\mathrm{GL}_{n}(K)$ be a representation, where $K$ is any field. Assume that for every fiber $F$ of $f$, the image $\varrho({\rm Im}[\pi_{1}(F)\to\pi_{1}(X)])$ is finite. Then there exists a finite étale covering $\mu:X^{\prime}\to X$ with the following property : Let $X^{\prime}\to S^{\prime}\to S$ be the Stein factorization of the composite $X^{\prime}\to X\to S$. Then for every fiber $G$ of $g:X^{\prime}\to S^{\prime}$, the image $\mu^{*}\varrho({\rm Im}[\pi_{1}(G)\to\pi_{1}(X^{\prime})])$ is trivial.
Démonstration.
We take $s\in S$. Since $\varrho(\pi_{1}(X))$ is residually finite and
$\varrho({\rm Im}[\pi_{1}(f^{-1}(s))\to\pi_{1}(X)])$
is finite, we may take a finite indexed normal subgroup $N_{s}\subset\varrho(\pi_{1}(X))$ such that
$\varrho({\rm Im}[\pi_{1}(f^{-1}(s))\to\pi_{1}(X)])\cap N_{s}=\{1\}.$
We take a connected open neighborhood $U_{s}\subset X$ of $f^{-1}(s)$ as in Lemma˜6.4. Then
$\varrho({\rm Im}[\pi_{1}(U_{s})\to\pi_{1}(X)])\cap N_{s}=\{1\}.$
We note that $f:X\to S$ is proper. Hence, we may take an open neighborhood $V_{s}\subset S$ of $s\in S$ such that $f^{-1}(V_{s})\subset U_{s}$. Then for $t\in V_{s}$, we have $f^{-1}(t)\subset U$, hence
$\varrho({\rm Im}[\pi_{1}(f^{-1}(t))\to\pi_{1}(X)])\cap N_{s}=\{1\}.$
Since $S$ is compact, we may take finite points $s_{1},\ldots,s_{k}\in S$ such that $S=V_{s_{1}}\cup\cdots\cup V_{s_{k}}$. Set $N=N_{s_{1}}\cap\cdots\cap N_{s_{k}}$. Then for every fiber $F$ of $f$, we have $\varrho({\rm Im}[\pi_{1}(F)\to\pi_{1}(X)])\cap N=\{1\}$.
Let $\mu:X^{\prime}\to X$ be a Galois covering corresponding to $\varrho^{-1}(N)\subset\pi_{1}(X)$. Let $X^{\prime}\to S^{\prime}\to S$ be the Stein factorization of the composite $X^{\prime}\to X\to S$. Then for every fiber $G$ of $g:X^{\prime}\to S^{\prime}$, we have ${\rm Im}[\pi_{1}(G)\to\pi_{1}(X)]\subset{\rm Im}[\pi_{1}(F)\to\pi_{1}(X)]$, where $F$ is the fiber of $f$ such that $\mu(G)=F$. Hence $\varrho({\rm Im}[\pi_{1}(G)\to\pi_{1}(X)])\cap N=\{1\}$. By $N=\varrho(\pi_{1}(X^{\prime}))$, we conclude $\mu^{*}\varrho({\rm Im}[\pi_{1}(G)\to\pi_{1}(X^{\prime})])=\{1\}$. ∎
Proposition 6.8.
[…]
Démonstration.
By Theorem˜2.9, the Shafarevich morphism ${\rm sh}_{\varrho}:X\to{\rm Sh}_{\varrho}(X)$ exists, and for each fiber $F$ of ${\rm sh}_{\varrho}$, $\varrho({\rm Im}[\pi_{1}(F)\to\pi_{1}(X)])$ is finite. By virtue of Lemma˜6.7, there exists a finite étale cover $\mu:\widehat{X}\to X$ such that considering the Stein factorization $\widehat{X}\stackrel{{\scriptstyle g}}{{\to}}S\to{\rm Sh}_{\varrho}(X)$ of the composite ${\rm sh}_{\varrho}\circ\mu$, for every fiber $G$ of $g$, the image $)\mu^{*}\varrho({\rm Im}[\pi_{1}(G)\to\pi_{1}(\widehat{X}$)]) is trivial. Applying Lemma˜6.6, we conclude that there exists a representation $\tau:\pi_{1}(S)\to{\rm GL}_{N}(K)$ such that $g^{*}\tau=\mu^{*}\varrho$.
We note that for each fiber $G$ of $g$, $g(G)$ is a point if and only if
$)\mu^{*}\varrho({\rm Im}[\pi_{1}(G^{\rm norm})\to\pi_{1}(\widehat{X}$)])
[…]
Let us prove that $\tau$ is large. Let $Z\subset S$ be a positive dimensional closed subvariety. Let $W\subset g^{-1}(Z)$ be an irreducible component that is dominant over $Z$. It follows that $)\mu^{*}\varrho({\rm Im}[\pi_{1}(W^{\rm norm})\to\pi_{1}(\widehat{X}$)]) is infinite by Theorem˜2.9. Since $g^{*}\tau=\mu^{*}\varrho$, it follows that
$\mu^{*}\varrho({\rm Im}[\pi_{1}(W^{\rm norm})\to\pi_{1}(\widehat{X})])\subset\tau({\rm Im}[\pi_{1}(Z^{\rm norm})\to\pi_{1}(S)])$
is infinite. This implies that $\tau$ is a large representation. ∎
In the rest of this paper, we will prove Theorem˜B.
6.3 Partial Albanese morphism
For this subsection we refer the readers to [Eys04, Eys11] for details. Let $X$ be a smooth projective variety. Let $\{\eta_{1},\ldots,\eta_{k}\}\subset H^{0}(X,\Omega_{X}^{1})$ be a set of holomorphic one forms. Consider the Albanese map ${\rm alb}_{X}:X\to{\rm Alb}(X)$. Then there exist holomorphic one forms $\{\omega_{1},\ldots,\omega_{k}\}\subset H^{0}({\rm Alb}(X),\Omega_{{\rm Alb}(X)}^{1})$ such that ${\rm alb}_{X}^{*}\omega_{i}=\eta_{i}$ for each $i$. Let $B$ be the largest abelian subvariety in ${\rm Alb}(X)$ such that $\omega_{i}|_{B}\equiv 0$ for each $i$. Denote by $A:={\rm Alb}(X)/B$ the quotient which is also an abelian variety. Then the morphism $a:X\to A$ that is the composite of ${\rm alb}_{X}$ and the quotient map $q:{\rm Alb}(X)\to A$ is called the partial Albanese morphism induced by $\{\eta_{1},\ldots,\eta_{k}\}$.
We remark that there exist holomorphic 1-forms $\{\omega^{\prime}_{1},\ldots,\omega^{\prime}_{k}\}\subset H^{0}(A,\Omega_{A}^{1})$ such that $q^{*}\omega_{i}^{\prime}=\omega_{i}$ for each $i$. They satisfy the following property :
Lemma 6.9.
For any positive-dimensional closed subvariety $Z$ of $a(X)$, there exists some $\omega_{i}^{\prime}$ such that $\omega_{i}^{\prime}|_{Z}\not\equiv 0$.
Démonstration.
Since $a:X\to a(X)$ is surjective, there exists an irreducible component $W$ in $a^{-1}(Z)$ that is dominant over $Z$. By the property of the partial Albanese map proved in [CDY22, Lemma 1.1], there exists some $\omega_{i}$ such that $\omega_{i}|_{W}\not\equiv 0$. Since $\omega_{i}=a^{*}\omega^{\prime}_{i}$, it follows that $\omega_{i}^{\prime}|_{Z}\not\equiv 0$. ∎
6.4 A recollection of spectral covering and canonical currents
Let $X$ be a smooth projective variety. Let $\tau:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be a reductive representation where $K$ is a non-archimedean local field. According to Theorem˜1.2, the Katzarkov-Eyssidieux reduction map $s_{\tau}:X\to S_{\tau}$ of $\tau$ exists, fulfilling the properties outlined therein. Since $X$ is projective, $s_{\tau}:X\to S_{\tau}$ is unique. We will outline the construction of certain canonical positive closed $(1,1)$-currents over $S_{\tau}$.
We first recall some facts about spectral forms as we have already seen in Item˜(a) in the proof of Theorem˜3.1. We refer the readers to [Eys04, CDY22, DYK23] for more details. Let
$\tau:\pi_{1}(X)\to{\rm GL}_{N}(K)$
be a reductive representation where $K$ is a non-archimedean local field. By the work of Gromov-Schoen [GS92] (see also [BDDM22] in cases where $X$ is non-compact), there exists a $\tau$-equivariant harmonic mapping $u:\widetilde{X}\to\Delta(G)$ where $\Delta(G)$ is the (enlarged) Bruhat-Tits building of $G$ (see [KP23, Definition 4.3.2] for the definition). Such a harmonic map is pluriharmonic, and the $(1,0)$-part of the complexified differentials gives rise to a set of multivalued holomorphic 1-forms over a dense open set of $X$ whose complement has Hausdorff codimension at least two. We can take some finite (ramified) Galois covering $\pi:X^{\!\rm sp}\to X$ (so-called spectral covering) with Galois group $H$ and such that the pullback of these multivalued one forms becomes single valued ones $\{\eta_{1},\ldots,\eta_{k}\}\subset H^{0}(X^{\!\rm sp},\pi^{*}\Omega_{X}^{1})$, that are called the spectral 1-forms of $\tau$. Consequently, the Stein factorization of the partial Albanese morphism $a:X^{\!\rm sp}\to A$ of $\{\eta_{1},\ldots,\eta_{k}\}$ leads to the (Katzarkov-Eyssidieux) reduction map $s_{\pi^{*}\tau}:X^{\!\rm sp}\to S_{\pi^{*}\tau}$ of $\pi^{*}\tau$. This map $s_{\pi^{*}\tau}$ is $H$-equivariant and its quotient by $H$ gives rise to the reduction map $s_{\tau}:X\to S_{\tau}$ by $\tau$. More precisely, we have the following commutative diagram :
[…]
Here $\sigma_{\pi}$ is also a finite ramified Galois cover with Galois group $H$. Note that there are 1-forms $\{\eta_{1}^{\prime},\ldots,\eta_{m}^{\prime}\}\subset H^{0}(A,\Omega_{A}^{1})$ such that $a^{*}\eta_{i}^{\prime}=\eta_{i}$. We define a smooth positive closed $(1,1)$-form $T_{\pi^{*}\tau}:=b^{*}\sum_{i=1}^{m}i\eta_{i}^{\prime}\wedge\overline{\eta_{i}}^{\prime}$ on $S_{\pi^{*}\tau}$. Note that $T_{\pi^{*}\tau}$ is invariant under the Galois action $H$. Therefore, there is a positive closed $(1,1)$-current $T_{\tau}$ defined on $S_{\tau}$ with continuous local potential such that $\sigma_{\pi}^{*}T_{\tau}=T_{\pi^{*}\tau}$.
Definition 6.10 (Canonical current).
The closed positive $(1,1)$-current $T_{\tau}$ on $S_{\tau}$ is called the canonical current of $\tau$.
Lemma 6.11.
$\{T_{\tau}\}$ is strictly nef. Namely, for any irreducible curve $C\subset S_{\tau}$, we have $\{T_{\tau}\}\cdot C>0$.
Démonstration.
Let $C^{\prime}\subset\sigma_{\pi}^{-1}(C)$ be an irreducible component which is dominant over $C$. Consider its image $b(C^{\prime})$. By Lemma˜6.9, there exists some $\eta_{i}^{\prime}\in H^{0}(A,\Omega_{A}^{1})$ such that $\eta_{i}^{\prime}|_{b(C^{\prime})}\neq 0$. Hence $i\eta_{i}^{\prime}\wedge\overline{\eta^{\prime}_{i}}|_{b(C^{\prime})}$ is strictly positive at general points. Consequently, $\{T_{\tau}\}\cdot C>0$. ∎
The canonical current $T_{\varrho}$ will serve as a lower bound for the complex hessian of plurisubharmonic functions constructed by the method of harmonic mappings.
Proposition 6.12 ([Eys04, Proposition 3.3.6, Lemme 3.3.12]).
Let $X$ be a projective normal variety and let $\varrho:\pi_{1}(X)\to G(K)$ be a Zariski dense representation where $K$ is a non archimedean local field and $G$ is a reductive group. Let $x_{0}\in\Delta(G)$ be an arbitrary point. Let $u:\widetilde{X}\to\Delta(G)$ be the associated harmonic mapping, where $\widetilde{X}$ is the universal covering of $X$. The function $\phi:\widetilde{X}\rightarrow\mathbb{R}_{\geq 0}$ defined by
$\phi(x)=2d^{2}\big(u(x),u(x_{0})\big)$
satisfies the following properties :
  1. (a)
    $\phi$ descends to a function $\phi_{\varrho}$ on $\widetilde{X_{\varrho}}=\widetilde{X}/\operatorname{ker}\left(\varrho\right)$.
  2. (b)
    Let $\Sigma$ be a normal complex space and $r:\widetilde{X}_{\varrho}\to\Sigma$ a proper holomorphic fibration such that $s_{\varrho}\circ\pi:\widetilde{X}_{\varrho}\rightarrow S_{\varrho}$ factorizes via a morphism $\nu:\Sigma{\rightarrow}S_{\varrho}$. The function $\phi_{\varrho}$ is of the form $\phi_{\varrho}=\phi_{\varrho}^{\Sigma}\circ r$ with $\phi_{\varrho}^{\Sigma}$ being a continuous plurisubharmonic function on $\Sigma$ ;
  3. (c)
    ${\rm d}{\rm d}^{\rm c}\phi^{\Sigma}_{\varrho}\geq\nu^{*}T_{\varrho}$.∎
We require the following criterion for an infinite topological Galois covering of a compact complex normal space to being Stein.
Proposition 6.13 ([Eys04, Proposition 4.1.1]).
Let $S$ be a compact complex normal space and let $\nu:\Sigma\to S$ be some infinite topological Galois covering. Let $T$ be a closed positive $(1,1)$-current on $S$ with continuous potential such that $\{T\}$ is a Kähler class. Assume that there exists a continuous plurisubharmonic function $\phi:\Sigma\to\mathbb{R}_{\geq 0}$ such that ${\rm d}{\rm d}^{\rm c}\phi\geq\nu^{*}T$. Then $\Sigma$ is a Stein space. ∎
6.5 Some lemmas in algebraic groups
We start with the following lemma.
Lemma 6.14.
Let $\Gamma_{1}$ be a finite index subgroup of $\Gamma_{2}$. Let $G$ be a linear algebraic group over a field $K$. Assume that $\Gamma_{2}\subset G(K)$ is Zariski dense. Then for the Zariski closure $H$ of $\Gamma_{1}$ in $G$, we have $H^{o}=G^{o}$, where $H^{o}$ and $G^{o}$ are the identity components of $H$ and $G$ respectively.
Démonstration.
We may assume that $K$ is algebraically closed. Replacing $\Gamma_{1}$ and $\Gamma_{2}$ by some finite index subgroups, we may assume that their Zariski closures are $H^{o}$ and $G^{o}$ respectively. Let $x_{1}\Gamma_{1},\ldots,x_{k}\Gamma_{1}$ be left cosets of $\Gamma_{1}$ in $\Gamma_{2}$. It follows that $x_{i}\Gamma_{1}$ is contained in the Zariski closed subset $x_{i}H$. Hence $G^{o}/H^{o}$ is finite. Since $G^{o}$ and $H^{o}$ are both connected, it implies that $G^{o}=H^{o}$. The lemma is proved. ∎
The following lemma plays a crucial role in the proof of Theorem˜6.23. It is a variant of [DYK23, Lemma 4.8] and its proof is based on [CDY22, Lemma 5.3].
Lemma 6.15.
Let $G$ be a semisimple algebraic group over a non-archimedean local field $K$. Let $\Gamma\subset G(K)$ be a finitely generated subgroup which is Zariski dense in $G$. If its derived group $\mathcal{D}\Gamma$ is bounded, then $\Gamma$ is also bounded.
Démonstration.
We can replace $\Gamma$ by a finite index subgroup and assume that $G$ is connected. Since $G$ is semisimple, according to the decomposition theorem [Mil17, Theorem 21.51] there are finitely many almost $K$-simple normal subgroups $H_{1},\ldots,H_{k}$ of $G$, such that $H_{1}\times\cdots\times H_{k}\to G$ is a central isogeny. Hence each quotient $G_{i}:=G/H_{1}\cdots H_{i-1}H_{i+1}\cdots H_{k}$ is an almost $K$-simple algebraic group. Let $\Gamma_{i}$ be the image of $\Gamma$ under the homomorphism $q_{i}:G(K)\to G_{i}(K)$. Then $\Gamma_{i}$ is Zariski dense in $G_{i}$, and we have $\mathcal{D}\Gamma_{i}=q_{i}(\mathcal{D}\Gamma)$. Since $\mathcal{D}\Gamma$ is bounded, by [KP23, Fact 2.2.4], $\mathcal{D}\Gamma_{i}$ is also bounded.
To show that $\Gamma_{i}$ is bounded, we assume contrary that $\Gamma_{i}$ is unbounded. Since $\mathcal{D}\Gamma_{i}$ is bounded and $G_{i}$ is almost $K$-simple, we apply [CDY22, Lemma 5.3] to conclude that $\mathcal{D}\Gamma_{i}$ is finite. Let $\Gamma^{\prime}_{i}$ by a finite index subgroup of $\Gamma_{i}$ such that $\mathcal{D}\Gamma^{\prime}_{i}$ is trivial. Then $\Gamma^{\prime}_{i}$ is commutative. Hence the Zariski closure of $\Gamma_{i}^{\prime}\subset G_{i}(K)$ is commutative. Since $\Gamma_{i}$ is Zariski dense in $G_{i}$, by Lemma˜6.14 we conclude that the identity component $G^{\circ}_{i}$ of $G_{i}$ is commutative. This contradicts with that $G_{i}$ is almost $K$-simple. Hence $\Gamma_{i}$ is bounded for each $i$.
Note that the natural morphism $G\to G_{1}\times\cdots\times G_{k}$ is also an isogeny. As a result, $\Gamma$ is bounded. ∎
6.6 Holomorphic convexity of universal covering
We recall the following criterion for an infinite topological Galois covering of a compact complex normal space to be Stein. As it is well known, we omit the proof.
Lemma 6.16.
Let $X$ be a projective normal variety and let $a:X\to A$ be a finite morphism to an abelian variety $A$. Then the universal covering $\widetilde{X}$ of $X$ is a Stein space.∎
We recall the notion of character varieties from sections §˜2.2 and §˜2.3. In particular, we set $M_{\rm B}(X,N)_{\mathbb{F}_{p}}=M_{\rm B}(\pi_{1}(X),N)_{\mathbb{F}_{p}}$.
Theorem 6.17.
Let $X_{0}$ be a complex projective normal surface and let $p$ be a fixed prime and $N$ be a fixed positive integer. We assume that $M_{\rm B}(X_{0},N)_{\mathbb{F}_{p}}$ is large, meaning that for any positive dimensional closed subvariety $Z$ of $X_{0}$, there exists a linear representation $\varrho:\pi_{1}(X_{0})\to{\rm GL}_{N}(K)$ with ${\rm char}\,K=p$ such that $\varrho({\rm Im}[\pi_{1}(Z^{\rm norm})\to\pi_{1}(X_{0})])$ is infinite. Then the universal covering of $X_{0}$ is Stein.
Démonstration.
Let $\mu:X\to X_{0}$ be a desingularization. It induces a morphism
$\iota:M_{\rm B}(X_{0},N)_{\mathbb{F}_{p}}\hookrightarrow M_{\rm B}(X,N)_{\mathbb{F}_{p}}$
between affine $\mathbb{F}_{p}$-schemes of finite type which is a closed immersion. Let $M:=\iota(M_{\rm B}(X_{0},N)_{\mathbb{F}_{p}})$ which is a Zariski closed subset of $M_{\rm B}(X,N)_{\mathbb{F}_{p}}$. By Theorem˜2.7, the Shafarevich morphism $s_{M}:X\to S_{M}$ of $M$ exists.
Claim 6.18.
The Shafarevich morphism $s_{M}:X\to S_{M}$ coincides with $\mu:X\to X_{0}$.
Démonstration.
Let $Z$ be any closed subvariety of $X$ and let $W:=\mu(Z)$. If $s_{M}(Z)$ is a point, then for any representation $\sigma:\pi_{1}(X_{0})\to{\rm GL}_{N}(L)$, where $L$ is any field of characteristic $p$ and $\sigma]\in M_{\mathrm{B}}(X_{0},N)_{\mathbb{F}_{p}}(L)$, the image $\mu^{*}\sigma(\mathrm{Im}[\pi_{1}(Z^{\mathrm{norm}})\to\pi_{1}(X)])$ is finite. Therefore, $\sigma({\rm Im}[\pi_{1}(W^{\rm norm})\to\pi_{1}(X_{0})])$ is finite. Hence $W$ is a point since $M_{\rm B}(X_{0},N)_{\mathbb{F}_{p}}$ is assumed to be large.
On the other hand, assume that $\mu(Z)$ is a point. Let $\sigma:\pi_{1}(X)\to{\rm GL}_{N}(L)$ be any representation where $L$ is a field of characteristic $p$ such that $\sigma]\in M(L)$. Then there exists a representation $\sigma^{\prime}:\pi_{1}(X_{0})\to{\rm GL}_{N}(\bar{L})$ such that $\mu^{*}\sigma^{\prime}]=[\sigma]$. By Lemma˜2.1, for any subgroup $\Gamma\subset\pi_{1}(X)$, $\mu^{*}\sigma^{\prime}(\Gamma)$ is finite if and only if $\sigma(\Gamma)$ is finite. Since $\mu^{*}\sigma^{\prime}({\rm Im}[\pi_{1}(Z^{\rm norm})\to\pi_{1}(X)])$ is trivial, it follows that $\sigma({\rm Im}[\pi_{1}(Z^{\rm norm})\to\pi_{1}(X)])$ is finite. By the properties of $s_{M}$ proven in Theorem˜2.7, $s_{M}(Z)$ is a point. This proves that $s_{M}=\mu$. ∎
Let $H_{0}:=\cap_{\varrho}\ker\varrho$ where $\varrho:\pi_{1}(X_{0})\to{\rm GL}_{N}(L)$ ranges over all linear representations and $L$ is any field with ${\rm char}\,L=p$. Denote by $H:=\mu_{*}^{-1}H_{0}$, where $\mu_{*}:\pi_{1}(X)\to\pi_{1}(X_{0})$ is the induced morphism. Let $\pi_{H}:\widetilde{X}_{H}\to X$ be the Galois covering with the Galois group $\pi_{1}(X)/H$ and $\pi_{0}:\widetilde{X}_{0}\to X_{0}$ be that with $\pi_{1}(X_{0})/H_{0}$. Consequently, we have the following commutative diagram
[…]
where $f:\widetilde{X}_{H}\to\widetilde{X}_{0}$ is a holomorphic proper fibration.
Claim 6.19.
Let $\sigma:\pi_{1}(X)\to{\rm GL}_{N}(L)$ be a linear representation such that $\sigma]\in M(L)$, where $L$ is a field of characteristic $p$. Then $H\subset\mathrm{ker}(\sigma^{ss})$.
Démonstration.
There exists a linear representation $\sigma^{\prime}:\pi_{1}(X_{0})\to{\rm GL}_{N}(\bar{L})$ such that $\mu^{*}\sigma^{\prime}]=[\sigma]$. Then $(\mu^{*}\sigma^{\prime})^{ss}$ is conjugate to $\sigma^{ss}$. Hence $\mathrm{ker}(\mu^{*}\sigma^{\prime})\subset\mathrm{ker}((\mu^{*}\sigma^{\prime})^{ss})=\mathrm{ker}(\sigma^{ss})$. We have $H\subset\mu^{*}\mathrm{ker}(\sigma^{\prime})=\mathrm{ker}(\mu^{*}\sigma^{\prime})$. Thus $H\subset\mathrm{ker}(\sigma^{ss})$. ∎
Consider the affine $\mathbb{F}_{p}$-scheme $R_{\mathbb{F}_{p}}$ and the GIT quotient $\pi:R_{\mathbb{F}_{p}}\to M_{\rm B}(X,N)_{\mathbb{F}_{p}}$, as defined in §˜2.1. Define $R:=\pi^{-1}(M)$, which is an affine $\overline{\mathbb{F}_{p}}$-scheme of finite type. According to the structure of the Shafarevich morphism described in Theorem˜2.7 and Definition˜2.6, the Shafarevich morphism $s_{M}:X\to S_{M}$ of $M$ is obtained through the simultaneous Stein factorization of the reductions $\{s_{\tau}:X\to S_{\tau}\}_{[\tau]\in M(K)}$. Here $\tau:\pi_{1}(X)\to{\rm GL}_{N}(K)$ ranges over all reductive representations with $K$ a local field of characteristic $p$ such that $\tau]\in M(K)$ and $s_{\tau}:X\to S_{\tau}$ is the reduction map defined in Theorem˜1.2.
By Lemma˜1.6, there exists a reductive representation ${\bm{\tau}}:\pi_{1}(X)\to{\rm GL}_{N^{\prime}}(K^{\prime})$ with $K^{\prime}$ a local field of characteristic $p$ such that the Katzarkov-Eyssidieux reduction map $s_{{\bm{\tau}}}:X\to S_{{\bm{\tau}}}$ of $\bm{\tau}$} coincides with $s_{M}:X\to S_{M}$. Moreover by Lemma˜1.6 and Claim˜6.19, we have
$H\subset\mathrm{ker}({\bm{\tau}}).$ (6.5)
For the spectral one forms $\{\eta_{1},\ldots,\eta_{k}\}\subset H^{0}(X^{\!\rm sp},\pi^{*}\Omega_{X}^{1})$ associated with the representation ${\bm{\tau}}:\pi_{1}(X)\to{\rm GL}_{N^{\prime}}(K^{\prime})$, we define the rank as follows : if there exist $i$ and $j$ such that $\eta_{i}\wedge\eta_{j}\neq 0$, then we say that the rank is $2$ ; otherwise, the rank is $1$.
Case 1 : The spectral 1-forms have rank 2. Assume that the spectral one forms
$\{\eta_{1},\ldots,\eta_{k}\}\subset H^{0}(X^{\!\rm sp},\pi^{*}\Omega_{X}^{1})$
with respect to ${\bm{\tau}}:\pi_{1}(X)\to{\rm GL}_{N^{\prime}}(K^{\prime})$ have rank 2.
Let $s_{\bm{\tau}}:X\to S_{\bm{\tau}}$ be the reduction map of $\bm{\tau}$}, which coincides with $s_{M}:X\to S_{M}$ (and thus $\mu:X\to X_{0}$). Let $T_{\bm{\tau}}$ be the canonical current on $X_{0}=S_{\bm{\tau}}$ defined in Definition˜6.10. Then $T_{\bm{\tau}}$ is strictly positive at general points since the spectral forms associated with $\bm{\tau}$} has rank 2. Since $T_{\bm{\tau}}$ has continuous local potentials, it follows that $T_{\bm{\tau}}$ is a big and nef class.
On the other hand, by Lemma˜6.11, we conclude that $T_{\bm{\tau}}$ is strictly nef. We now apply a theorem by Demailly-Păun [DP04] (or [Lam99]) to conclude that $\{T_{\bm{\tau}}\}$ is a Kähler class.
According to Proposition˜6.12 and (6.5), there exists a continuous plurisubharmonic function $\phi:\widetilde{X}_{0}\to\mathbb{R}_{\geq 0}$ such that ${\rm d}{\rm d}^{\rm c}\phi\geq\pi_{0}^{*}T_{{\bm{\tau}}}$. We can apply Proposition˜6.13 to conclude that $\widetilde{X}_{0}$ is a Stein space. Note that any unramified covering of a Stein space is Stein. Therefore, the universal covering of $X_{0}$ is a Stein space.
Case 2 : The spectral 1-forms have rank 1 : Assume that the spectral covering $X^{\!\rm sp}\to X$ with respect to ${\bm{\tau}}:\pi_{1}(X)\to{\rm GL}_{N^{\prime}}(K^{\prime})$ has rank 1 ; i.e., $\eta_{i}\wedge\eta_{j}\equiv 0$ for every $\eta_{i}$ and $\eta_{j}$, where $\{\eta_{1},\ldots,\eta_{\ell}\}\subset H^{0}(X^{\!\rm sp},\pi^{*}\Omega_{X}^{1})$ is the spectral forms associated with ${\bm{\tau}}:\pi_{1}(X)\to{\rm GL}_{N^{\prime}}(K^{\prime})$, and $\pi:X^{\!\rm sp}\to X$ is the spectral covering associated with ${\bm{\tau}}:\pi_{1}(X)\to{\rm GL}_{N^{\prime}}(K^{\prime})$.
Case 2.1 : The dimension of spectral 1-forms is at least two : Suppose ${\rm dim}_{\mathbb{C}}{\rm Span}\{\eta_{1},\ldots,\eta_{\ell}\}\geq 2.$
Without loss of generality, we may assume that $\eta_{1}\wedge\eta_{2}\equiv 0$ and $\eta_{1}\not\in\{\mathbb{C}\eta_{2}\}$. According to the Castelnuovo-De Franchis theorem (cf. [ABC+96, Theorem 2.7]), there exists a proper fibration $h:X^{\!\rm sp}\to C$ over a smooth projective curve $C$ such that $\{\eta_{1},\eta_{2}\}\subset h^{*}H^{0}(C,\Omega_{C}^{1})$. Since $s_{M}=\mu$ is birational, we can choose a general fiber $F$ of $h$, which is irreducible and such that $s_{M}\circ\pi(F)=s_{\bm{\tau}}\circ\pi(F)$ is not a point. There exists some $i$ such that $\eta_{i}|_{F}\neq 0$ (cf. Lemma˜6.9). Given that $\eta_{1}|_{F}\equiv 0$, this implies that $\eta_{i}\wedge\eta_{1}\neq 0$. It contradicts with our assumption that the spectral $1$-forms have rank 1. Therefore, this case cannot occur.
Case 2.2 : The dimension of spectral 1-forms is 1 : We have ${\rm dim}_{\mathbb{C}}{\rm Span}\{\eta_{1},\ldots,\eta_{\ell}\}=1.$
Let $G$ be the Zariski closure of ${\bm{\tau}}(\pi_{1}(X))$, which is reductive. Consider the isogeny $g:G\to G/Z\times G/\mathcal{D}G$ where $Z$ is the central torus of $G$ and $\mathcal{D}G$ is the derived group of $G$. As a result, $G^{\prime}:=G/Z$ is semisimple and $G^{\prime\prime}:=G/\mathcal{D}G$ is a torus. Let $\tau^{\prime}:\pi_{1}(X)\to G^{\prime}(\overline{K^{\prime}})$ be the composite of $\bm{\tau}$} with the projection $G\to G^{\prime}$, and $\tau^{\prime\prime}:\pi_{1}(X)\to G^{\prime\prime}(\overline{K^{\prime}})$ be the composite of $\bm{\tau}$} with the projection $G\to G^{\prime\prime}$. Then $\tau^{\prime}$ and $\tau^{\prime\prime}$ are both Zariski dense representations.
Let $\nu:Y\to X^{\!\rm sp}$ be a desingularization and denote $\eta:=\nu^{*}\eta_{1}$. Consider the partial Albanese morphism $a:Y\to A$ induced by $\eta$. Then there exists a one form $\eta^{\prime}\in H^{0}(A,\Omega_{A}^{1})$ such that $a^{*}\eta^{\prime}=\eta$. If $\dim a(Y)=1$, then the Stein factorization $h:Y\to C$ of $a$ is a proper holomorphic fibration over a smooth projective curve $C$ such that $\eta_{1}\in h^{*}H^{0}(C,\Omega_{C}^{1})$. We are now in a situation akin to Case 2.1, and we can apply the same arguments to reach a contradiction. Hence $\dim a(Y)=2$. Let $\pi_{A}:\widetilde{A}\to A$ denote the universal covering map. We denote by $Y^{\prime}:=Y\times_{\widetilde{A}}A$ a connected component of the fiber product and let $\pi^{\prime}:Y^{\prime}\to Y$ be the induced étale cover. It’s worth noting that $\pi^{\prime*}\eta$ is exact. Consequently, we can define the following holomorphic map :
$h:Y^{\prime}$ $\to\mathbb{C}$
$y$ $\mapsto\int_{y_{0}}^{y}\pi^{\prime*}\eta.$
We then have the following commutative diagram :
[…]
The holomorphic map $\widetilde{A}\to\mathbb{C}$ in the above diagram is defined by the linear 1-form $\pi_{A}^{*}\eta^{\prime}$ on $\widetilde{A}$. By Simpson’s Lefschetz theorem [Sim93a], for any $t\in\mathbb{C}$, $h^{-1}(t)$ is connected and $\pi_{1}(h^{-1}(t))\to\pi_{1}(Y^{\prime})$ is surjective. By definition of $h$, $\pi_{Y}^{*}\eta|_{Z}\equiv 0$ where $Z$ is any connected component of $p^{-1}(h^{-1}(t))$. Here $p:\widetilde{Y}\to Y^{\prime}$ is the natural covering map.
Consider the Zariski dense representation $\tau^{\prime}:\pi_{1}(X)\to G^{\prime}\big(\overline{K^{\prime}}\big)$ as defined previously. Let $L$ be a finite extension of $K^{\prime}$ such that $G^{\prime}$ is defined on $L$ and $\tau^{\prime}:\pi_{1}(X)\to G^{\prime}(L)$. We denote by $\sigma:\pi_{1}(Y)\to G^{\prime}\big(L\big)$ the pullback of $\tau^{\prime}$ via the morphism $Y\to X$. The existence of a $\sigma$-equivariant harmonic mapping $u:\widetilde{Y}\to\Delta(G^{\prime})$ is guaranteed by [GS92], where $\Delta(G^{\prime})$ is the Bruhat-Tits building of $G^{\prime}$.
We note that $\pi_{Y}^{*}\eta$ is the (1,0)-part of the complexified differential of the harmonic mapping $u$ at general points of $\widetilde{Y}$, with $\pi_{Y}:\widetilde{Y}\to Y$ denoting the universal covering. For any connected component $Z$ of $p^{-1}(h^{-1}(t))$ for a general $t\in\mathbb{C}$, since $\pi_{Y}^{*}\eta|_{Z}\equiv 0$, and all the spectral forms are assumed to be $\mathbb{C}$-linearly equivalent, it follows that $u(Z)$ is constant. Since $u$ is $\sigma$-equivariant, it follows that $\pi^{\prime*}\sigma({\rm Im}[\pi_{1}(h^{-1}(t))\to\pi_{1}(Y^{\prime})])$ is contained in the subgroup of $G^{\prime}(L)$ fixing the point $u(Z)$. Recall that $\pi_{1}(h^{-1}(t))\to\pi_{1}(Y^{\prime})$ is surjective. Hence $\pi^{\prime*}\sigma(\pi_{1}(Y^{\prime}))$ is a bounded subgroup of $G^{\prime}(L)$. Additionally, note that
$\mathcal{D}\pi_{1}(Y)\subset{\rm Im}[\pi_{1}(Y^{\prime})\to\pi_{1}(Y)],$
and it follows that $\sigma(\mathcal{D}\pi_{1}(Y))$ is bounded. Since $\tau^{\prime}$ is Zariski dense, and ${\rm Im}[\pi_{1}(Y)\to\pi_{1}(X)]$ is a finite index subgroup of $\pi_{1}(X)$, according to Lemma˜6.14 the Zariski closure of $\sigma(\pi_{1}(Y))$ contains the identity component of $G^{\prime}$, and it is also semisimple. We apply Lemma˜6.15 to conclude that $\sigma(\pi_{1}(Y))$ is bounded.
Since $\sigma(\pi_{1}(Y))$ is a finite index subgroup of $\tau^{\prime}(\pi_{1}(X))$, it follows that $\tau^{\prime}(\pi_{1}(X))$ is also bounded. Then the reduction map $s_{\tau^{\prime}}$ is the constant map. This implies that the reduction map $s_{\tau}$ is identified with $s_{\tau^{\prime\prime}}$. Recall that $G^{\prime\prime}$ is a tori. By [CDY22, Step 6 in Proof of Theorem H], we know that there exists a morphism $a_{X}:X\to A_{X}$ with $A_{X}$ an abelian variety such that $s_{\tau^{\prime\prime}}$ is the Stein factorization of $a_{X}$.
Claim 6.20.
There exists a morphism $b:X_{0}\to A_{X}$ such that $b\circ\mu=a_{X}$.
Démonstration.
Recall that $\tau]\in M(K^{\prime})$. Let $F$ be any fiber of the birational morphism $\mu:X\to X_{0}$. Then $\tau({\rm Im}[\pi_{1}(F)\to\pi_{1}(X)])$ is finite since $s_{M}:X\to S_{M}$ coincides with $\mu$. It follows that $\tau({\rm Im}[\pi_{1}(F)\to\pi_{1}(X)])$ is a finite group. By the definition of $\tau^{\prime\prime}$, we conclude that $\tau^{\prime\prime}({\rm Im}[\pi_{1}(F)\to\pi_{1}(X)])$ is also finite, and is thus bounded. According to the property of the reduction map Theorem˜1.2, $s_{\tau^{\prime\prime}}(F)$ is a point. Hence there exists a morphism $b:X_{0}\to A$ such that $b\circ\mu=a_{X}$. ∎
Consider the map $b:X_{0}\to A_{X}$. Note that this map is finite. We apply Lemma˜6.16 to conclude that the universal covering of $X_{0}$ is a Stein space. ∎
Remark 6.21.
The proof of Theorem˜6.17 utilizes techniques similar to those used in the proof of [DYK23, Theorem C]. In order to extend Theorem˜6.17 to any projective normal variety, we have to establish Simpson’s theory [Sim93b] on absolutely constructible subsets for character varieties of representations in positive characteristic. We plan to explore this problem in our future work.
We recall the following definition by Campana [Cam94].
Definition 6.22 ($\Gamma$-dimension).
Let $X$ be a projective normal variety. The $\Gamma$-dimension of $X$ is defined to be $\dim{\rm Sh}(X)$, where ${\rm sh}_{X}:X\dashrightarrow{\rm Sh}(X)$ is the Shafarevich map constructed by Campana [Cam94] and Kollár [Kol93].
Theorem 6.23.
Let $X$ be a projective normal variety and let $\varrho:\pi_{1}(X)\to{\rm GL}_{N}(K)$ be a faithful representation where $K$ is a field of positive characteristic. If the $\Gamma$-dimension of $X$ is at most two (e.g. when $\dim X\leq 2$), then the universal covering $\widetilde{X}$ of $X$ is holomorphically convex.
Démonstration.
By Proposition˜6.8, after we replace $X$ by a suitable finite étale cover, there exists a large representation $\tau:\pi_{1}({\rm Sh}_{\varrho}({X}))\to{\rm GL}_{N}(K)$ such that $({\rm sh}_{\varrho})^{*}\tau=\varrho$, where ${\rm sh}_{\varrho}:{X}\to{\rm Sh}_{\varrho}({X})$ is the Shafarevich morphism of $\varrho$. Since $\varrho$ is faithful, it follows that $\tau$ is also faithful since the homomorphism $({\rm sh}_{\varrho})_{*}:\pi_{1}(X)\to\pi_{1}({\rm sh}_{\varrho}(X))$ is surjective. As we assume that the $\Gamma$-dimension of $X$ is at most two, therefore, $\dim{\rm Sh}_{\varrho}({X})\leq 2$. We apply Theorem˜6.17 to conclude that the universal covering $S$ of ${\rm Sh}_{\varrho}({X})$ is Stein. Note that $({\rm sh}_{\varrho})_{*}:\pi_{1}(X)\to\pi_{1}({\rm sh}_{\varrho}(X))$ is an isomorphism. Hence there exists a proper holomorphic fibration $\widetilde{X}\to S$ between the universal coverings of $X$ and ${\rm Sh}_{\varrho}(X)$ that lifts ${\rm sh}_{\varrho}$. It follows that $\widetilde{X}$ is holomorphically convex. ∎
Acknowledgment.
We would like to thank Michel Brion, Benoît Claudon, Philippe Eyssidieux and Andreas Höring for very helpful discussions. We also thank Benoît Cadorel and Yuan Liu for reading the paper and their helpful remarks. Finally, we sincerely thank the referees for very careful readings and helpful remarks, in particular for pointing out a simpler proof of Corollary˜G (see Remark˜6.3). Y.D. acknowledges support from the ANR grant Karmapolis (ANR-21-CE40-0010). K.Y. acknowledges support from JSPS Grant-in-Aid for Scientific Research (C) 22K03286.
Références
  • [ABC+96] J. Amorós, M. Burger, K. Corlette, D. Kotschick & D. Toledo. Fundamental groups of compact Kähler manifolds, vol. 44 of Mathematical Surveys and Monographs. American Mathematical Society, Providence, RI (1996). URL http://dx.doi.org/10.1090/surv/044. $\uparrow$
  • [BDDM22] D. Brotbek, G. Daskalopoulos, Y. Deng & C. Mese. “Pluriharmonic maps into buildings and symmetric differentials”. arXiv e-prints, (2022) :arXiv :2206.11835. 2206.11835, URL http://dx.doi.org/10.48550/arXiv.2206.11835. $\uparrow$
  • [Bru23] Y. Brunebarbe. “Existence of the Shafarevich morphism for semisimple local systems on quasi-projective varieties”. arXiv e-prints, (2023) :arXiv :2305.09741. 2305.09741, URL http://dx.doi.org/10.48550/arXiv.2305.09741. $\uparrow$
  • [Cam94] F. Campana. “Remarks on the universal covering of compact Kähler manifolds”. Bull. Soc. Math. Fr., 122(1994)(2) :255–284. URL %****␣Shafarevich_morphism_final_version.tex␣Line␣2775␣****http://dx.doi.org/10.24033/bsmf.2232. $\uparrow$
  • [Cam04] ———. “Orbifolds, special varieties and classification theory.” Ann. Inst. Fourier, 54(2004)(3) :499–630. URL http://dx.doi.org/10.5802/aif.2027. $\uparrow$
  • [Cam11] ———. “Special orbifolds and birational classification : a survey”. “Classification of algebraic varieties. Based on the conference on classification of varieties, Schiermonnikoog, Netherlands, May 2009.”, 123–170. Zürich : European Mathematical Society (EMS) (2011) :. $\uparrow$
  • [CCE15] F. Campana, B. Claudon & P. Eyssidieux. “Linear representations of Kähler groups : factorizations and linear Shafarevich conjecture”. Compos. Math., 151(2015)(2) :351–376. URL http://dx.doi.org/10.1112/S0010437X14007751. $\uparrow$
  • [CDY22] B. Cadorel, Y. Deng & K. Yamanoi. “Hyperbolicity and fundamental groups of complex quasi-projective varieties”. arXiv e-prints, (2022) :arXiv :2212.12225. 2212.12225, URL http://dx.doi.org/10.48550/arXiv.2212.12225. $\uparrow$
  • [CH10] B. Claudon & A. Hoering. “Threefolds with quasi-projective universal cover”. arXiv e-prints, (2010) :arXiv :1009.3739. 1009.3739, URL http://dx.doi.org/10.48550/arXiv.1009.3739. $\uparrow$
  • [CH13] B. Claudon & A. Höring. “Compact Kähler manifolds with compactifiable universal cover”. Bull. Soc. Math. France, 141(2013)(2) :355–375. With an appendix by Frédéric Campana, URL http://dx.doi.org/10.24033/bsmf.2651. $\uparrow$
  • [CHK13] B. Claudon, A. Höring & J. Kollár. “Algebraic varieties with quasi-projective universal cover”. J. Reine Angew. Math., 679(2013) :207–221. URL http://dx.doi.org/10.1515/crelle.2012.017. $\uparrow$
  • [Con14] B. Conrad. “Reductive group schemes”. “Autour des schémas en groupes. Vol. I”, vol. 42/43 of Panor. Synthèses, 93–444. Soc. Math. France, Paris (2014) :. $\uparrow$
  • [CP19] F. Campana & M. Păun. “Foliations with positive slopes and birational stability of orbifold cotangent bundles”. Publ. Math., Inst. Hautes Étud. Sci., 129(2019) :1–49. URL http://dx.doi.org/10.1007/s10240-019-00105-w. $\uparrow$
  • [CW16] F. Campana & J. Winkelmann. “Rational connectedness and order of non-degenerate meromorphic maps from $\mathbb{C}^{n}$”. Eur. J. Math., 2(2016)(1) :87–95. URL http://dx.doi.org/10.1007/s40879-015-0083-z. $\uparrow$
  • [DP04] J.-P. Demailly & M. Paun. “Numerical characterization of the Kähler cone of a compact Kähler manifold”. Ann. Math. (2), 159(2004)(3) :1247–1274. URL http://dx.doi.org/10.4007/annals.2004.159.1247. $\uparrow$
  • [DYK23] Y. Deng, K. Yamanoi & L. Katzarkov. “Reductive Shafarevich Conjecture”. arXiv e-prints, (2023) :arXiv :2306.03070. 2306.03070, URL http://dx.doi.org/10.48550/arXiv.2306.03070. $\uparrow$
  • [EKPR12] P. Eyssidieux, L. Katzarkov, T. Pantev & M. Ramachandran. “Linear Shafarevich conjecture”. Ann. Math. (2), 176(2012)(3) :1545–1581. URL http://dx.doi.org/10.4007/annals.2012.176.3.4. $\uparrow$
  • [Eys04] P. Eyssidieux. “Sur la convexité holomorphe des revêtements linéaires réductifs d’une variété projective algébrique complexe”. Invent. Math., 156(2004)(3) :503–564. URL http://dx.doi.org/10.1007/s00222-003-0345-0. $\uparrow$
  • [Eys11] P. Eyssidieux. “Lectures on the Shafarevich conjecture on uniformization”. “Complex manifolds, foliations and uniformization. Collected papers of the summer school “Uniformisation de familles de variétés complexes”, Dijon, France, August 31 – September 11, 2009”, 101–148. Paris : Société Mathématique de France (SMF) (2011) :. $\uparrow$
  • [Fuj17] O. Fujino. “Notes on the weak positivity theorems”. “Algebraic varieties and automorphism groups. Proceedings of the workshop held at RIMS, Kyoto University, Kyoto, Japan, July 7–11, 2014”, 73–118. Tokyo : Mathematical Society of Japan (MSJ) (2017) :. $\uparrow$
  • [GS92] M. Gromov & R. Schoen. “Harmonic maps into singular spaces and $p$-adic superrigidity for lattices in groups of rank one”. Publ. Math., Inst. Hautes Étud. Sci., 76(1992) :165–246. URL %****␣Shafarevich_morphism_final_version.tex␣Line␣2925␣****http://dx.doi.org/10.1007/BF02699433. $\uparrow$
  • [Hof09] K. R. Hofmann. “Triangulation of locally semi-algebraic spaces.” https://deepblue.lib.umich.edu/bitstream/handle/2027.42/63851/krhofman_1.pdf ?sequence=1&isAllowed=y (2009). $\uparrow$
  • [Kat97] L. Katzarkov. “On the Shafarevich maps”. “Algebraic geometry. Proceedings of the Summer Research Institute, Santa Cruz, CA, USA, July 9–29, 1995”, 173–216. Providence, RI : American Mathematical Society (1997) :. $\uparrow$
  • [Kol93] J. Kollár. “Shafarevich maps and plurigenera of algebraic varieties”. Invent. Math., 113(1993)(1) :177–215. URL http://dx.doi.org/10.1007/BF01244307. $\uparrow$
  • [Kol95] ———. Shafarevich maps and automorphic forms. Princeton University Press, Princeton (N.J.) (1995). $\uparrow$
  • [KP23] T. Kaletha & G. Prasad. Bruhat-Tits theory. A new approach (to appear), vol. 44 of New Math. Monogr. Cambridge : Cambridge University Press (2023). $\uparrow$
  • [KR98] L. Katzarkov & M. Ramachandran. “On the universal coverings of algebraic surfaces”. Ann. Sci. École Norm. Sup. (4), 31(1998)(4) :525–535. URL http://dx.doi.org/10.1016/S0012-9593(98)80105-5. $\uparrow$
  • [Lam99] A. Lamari. “The Kähler cone of a surface”. J. Math. Pures Appl. (9), 78(1999)(3) :249–263. URL http://dx.doi.org/10.1016/S0021-7824(98)00005-1. $\uparrow$
  • [LM85] A. Lubotzky & A. R. Magid. “Varieties of representations of finitely generated groups”. Mem. Amer. Math. Soc., 58(1985)(336) :xi+117. URL http://dx.doi.org/10.1090/memo/0336. $\uparrow$
  • [Mil17] J. S. Milne. Algebraic groups. The theory of group schemes of finite type over a field, vol. 170 of Camb. Stud. Adv. Math. Cambridge : Cambridge University Press (2017). URL http://dx.doi.org/10.1017/9781316711736. $\uparrow$
  • [NWY13] J. Noguchi, J. Winkelmann & K. Yamanoi. “Degeneracy of holomorphic curves into algebraic varieties. II”. Vietnam J. Math., 41(2013)(4) :519–525. URL http://dx.doi.org/10.1007/s10013-013-0051-1. $\uparrow$
  • [Ses77] C. S. Seshadri. “Geometric reductivity over arbitrary base”. Advances in Math., 26(1977)(3) :225–274. URL http://dx.doi.org/10.1016/0001-8708(77)90041-X. $\uparrow$
  • [Sim88] C. T. Simpson. “Constructing variations of Hodge structure using Yang-Mills theory and applications to uniformization”. J. Amer. Math. Soc., 1(1988)(4) :867–918. URL http://dx.doi.org/10.2307/1990994. $\uparrow$
  • [Sim92] ———. “Higgs bundles and local systems”. Inst. Hautes Études Sci. Publ. Math., (1992)(75) :5–95. URL http://www.numdam.org/item ?id=PMIHES_1992__75__5_0. $\uparrow$
  • [Sim93a] ———. “Lefschetz theorems for the integral leaves of a holomorphic one-form”. Compos. Math., 87(1993)(1) :99–113. $\uparrow$
  • [Sim93b] ———. “Subspaces of moduli spaces of rank one local systems”. Ann. Sci. Éc. Norm. Supér. (4), 26(1993)(3) :361–401. URL http://dx.doi.org/10.24033/asens.1675. $\uparrow$
  • [ST00] G. Smith & O. Tabachnikova. Topics in group theory. Springer Undergraduate Mathematics Series. Springer-Verlag London, Ltd., London (2000). URL http://dx.doi.org/10.1007/978-1-4471-0461-2. $\uparrow$
  • [Yam10] K. Yamanoi. “On fundamental groups of algebraic varieties and value distribution theory”. Ann. Inst. Fourier, 60(2010)(2) :551–563. URL http://dx.doi.org/10.5802/aif.2532. $\uparrow$